### v2

# results on dynamics of bungee set of composite entire functions in the eremenko-lyubich class

Abstract.
In this paper, we have discussed the dynamics of composite entire functions in terms of relationship between bungee set, repelling periodic points (to be denoted by $RP$) and rationally indifferent fixed point set. We have established relation between the dynamics of composite function and the functions taken for composition. We showed that the union of their Bungee sets contains Bungee set of the composite function. We also showed that $RP$ set of composite functions contains the $RP$ set of functions used for the composition. We have mostly dealt with the permutable functions of class $\mathcal{B}$. These results hold true when we have a composite function in which one function is a periodic translation of another.
Key words and phrases:
bungee set, wandering domain, Baker domain, repelling periodic point, rationally indifferent fixed point, escaping set, asymptotic value, singular set
2020 Mathematics Subject Classification:
37F10, 30D05
The research work of the first author is supported by ANRF(SERB) research grant TAR/2023/000197
1. Introduction
When we iterate a homomorphic function $f:\mathbb{C}\rightarrow\mathbb{C}$, $n-$times, we observe that the points of the complex plane move differently, which we call the dynamics of the point where we denote $f^{n}$ to be $n-$th iteration of a function, where $n\in\mathbb{N}$. As, not all points have same dynamics, we therefore have different classifications of points of the complex plane $\mathbb{C}$; one classification is based on the nature of neighbourhood of a point $z_{0}$ (say); they are said to be in the Fatou set $F(f)$ (set of points whose neighbourhood exhibits stable behaviour) and Julia set $J(f)$ (set of points whose neighbourhood exhibits chaotic behaviour). This was broadly the classification we generally see when we talk about dynamics of points in complex plane.
There is yet another classification of these points based on the nature of their orbits; viz., the Filled in Julia set $K(f)$ (set of points whose orbits are bounded), escaping set $I(f)$ (set of points whose orbits tend to infinity), and bungee set $BU(f)$ (the orbits of such points neither remain bounded nor tend to infinity; only there exists a subsequence that remains bounded and another one that is unbounded). These sets are disjoint among themselves and their union $BU(f)\cup I(f)\cup K(f)=\mathbb{C}$. In [16], it was shown that if a Fatou component intersects with $BU(f)$ then that component becomes a wandering domain contained in $BU(f)$. $K(f)$ comprises of Fatou components in which orbits under iteration stays bounded, along with repelling periodic points and rationally indifferent periodic points that lies inside Julia set $J(f).$ It is to be noted that the bungee set and escaping set must have a finite intersection with $J(f)$ (see [5, 16]). The escaping set too has a finite intersection with the Fatou set when the Fatou component is a Baker domain. Knowing all these sets for an entire function, we were curious about what happens when we compose two functions and how these sets relate to the composition with those of individual functions. We tried to see what happens to the dynamics of transcendental functions under compositions as composition of two transcendental functions is also transcendental.
Simplifying the situation further, we have taken transcendental functions $f,g$ of Eremenko Lyubich class $\mathcal{B}$ which means that the functions of this class have a bounded set of singular values (i.e., the set of all asymptotic values and critical values and their finite limit points is bounded). Additionally, this class does not have any intersection of the Fatou set with the bungee set, in other words there is no wandering domain which also lies inside the bungee set for $f\in\mathcal{B}.$ Since our aim was to establish a relationship between different sets of functions and their composition and it would be quite difficult to establish any relation where there is no algebraic relation among these functions, we have focused on commuting(permutable) functions, that is, $f\circ g=g\circ f$, and $g=f+C$ where $C$ is the period of $f$.
The following results are proven in this paper:
Let $f,g\in\mathcal{B}$ be commuting functions. The bungee set of composed functions is contained in the union of the bungee set of functions used for composition. The set of repelling periodic and indifferent fixed points of the composed function $RP(f\circ g)$ contains the union of repelling periodic and indifferent fixed points of these functions; that is, $RP(f\circ g)\supset RP(f)\cup RP(g)$.
We know some results for exponential functions which belong to Speiser class $\mathcal{S}$ (such class of functions has finite number of elements in their singular set) [18]. Also, $\mathcal{S}\subset\mathcal{B}$. $f=e^{\lambda z},\lambda\neq 0$, a function $g$ that commutes under composition arewith $f$ is some finite iteration of $f$ itself, that is, $g=f^{n}$ for some $n\in\mathbb{N}$ [2]; therefore, for such a case, all sets of composed functions are equal to the functions taken for composition.
1.1. Observations
While developing the relationships among the functions and the individual functions, we also found that the bungee set and escaping set exhibit similar characteristics for functions of class $\mathcal{B}$:
  1. (1)
    The closure of both these sets equals Julia set.
  2. (2)
    Both sets are neither open nor closed (3.2).
  3. (3)
    Both sets of composite function is contained in union of individual function (to be proved in this paper).
2. Main results
To begin with the commuting entire functions, we first understand the nature of these sets when functions have no asymptotic values in their singular set.
Theorem 2.1.
If $f$ and $g$ are commuting functions having no finite asymptotic values, then $I(f)$ is completely invariant under $g$ and vice versa.
Proof.
[…]
But assuming $g(z)$ does not have any finite asymptotic value, so $|g(f^{n}(z_{0})|\rightarrow\infty$.
Since $f\,\&\,g$ are commuting so, $|g(f^{n}(z_{0})|$ = $|f^{n}(g(z_{0})|\rightarrow\infty.$ This implies that $g(z)$ carries $I(f)$ inside itself or in other words, $I(f)$ is forward invariant under $g$. It is easy to see that $g^{-1}(I(f))\subset I(f)$. Hence, $I(f)$ is completely invariant under $g$. As $f\,\&\,g$ are commuting, their roles can be interchanged and hence, $I(g)$ is completely invariant under $f$ as well. Hence, $I(f)$ and $I(g)$ is completely invariant under $f\circ g$ as well. ∎
Theorem 2.2.
Suppose $f$ and $g$ are commuting entire functions having no asymptotic values. Then $K(f)$ and $BU(f)$ are completely invariant under $g$.
Proof.
$g(K(f))$ is, in general forward invariant, because if we assume a point $w\in K(f)$, then we have $|f^{n}(w)|<R$ for some $R.$ It is clear that $|f^{m_{k}}(g(w))|<R_{1}$ for some $R_{1}$ which implies that $g(w)\in K(f)$, thereby proving the result. Now, let $z_{0}\in g^{-1}(K(f)).$ This implis that $g(z_{0})\in K(f)$ and $|f^{n}(g(z_{0}))|<L$ for some $L$ and because of commutation, $|g(f^{n}(z_{0}))|<L.$ As these functions do not have any finite asymptotic values, $|f^{n}(z_{0})|<L_{1}$ for some $L_{1}$, which means $z_{0}\in K(f)$. Thus, we have that $K(f)$ is completely invariant under $g$ when the functions have no finite asymptotic values.
Now bungee set, $BU(f)=\mathbb{C}\backslash(I(f))\cup K(f)).$ We know that $I(f)$ and $K(f)$ are completely invariant under $g$ (2.1) and hence, $BU(f)$ is also completely invariant under $g$. ∎
Lemma 2.3.
The filled-in Julia set of the composite function $f\circ g$ cannot be larger than intersection of filled in Julia set of the individual functions having no finite asymptotic values.
Proof.
We assume that $K(f\circ g)\supset K(f)\cap K(g)$. Thus, there might be points in $K(f\circ g)$ that have a finite intersection with $I(f)\cup I(g)\cup BU(f)\cup BU(g)\cup A\cup B$ where $A:K(f)\backslash K(f)\cap K(g)$ and $B:K(g)\backslash K(f)\cap K(g)$. If $z_{0}\in I(f)\cup I(g)\cup BU(f)\cup BU(g)$, these points cannot be bounded as one function carries the points of other functions to itself, meaning $f^{n}\circ g^{n}(z_{0})$ as $n\rightarrow\infty$ will either remain oscillatory like bungee points nor will tend to $\infty$. Now, let $w_{0}\in A$ and apply $f^{m}\circ g^{2m}\circ f^{m}(w_{0})$ as $m\rightarrow\infty$, which is bounded because $g(K(f))\subset K(f)$. However, $g^{m}\circ f^{2m}\circ g^{m}(w_{0})$, which cannot be bounded ($f(BU(g)\subset BU(g)$ and $g^{m}\circ f^{2m}\circ g^{m}(w_{0})$ cannot be bounded. As these functions have no finite asymptotic value, so $g^{m}\circ f^{2m}\circ g^{m}(w_{0})$ will be unbounded in the case when $w_{0}\in I(g).$) Similarly, such arguments may be given when $w_{1}\in B$. Thus, we arrive at a contradiction. Hence, $K(f\circ g)\subset K(f)\cap K(g)$. ∎
Theorem 2.4.
Commuting transcendental entire functions with no finite asymptotic values have equal bungee sets, filled-in Julia sets and escaping sets.
Proof.
Let $z_{0}\in BU(f\circ g)$ where $f\ \&\ g$ do not have any finite asymptotic value. To begin with let there be two subsequences $\{m_{k}\}$ and $\{n_{k}\}$ both tending to $\infty$ such that $|(f\circ g)^{m_{k}}(z_{0})|<R$ for some $R$ and $|(f\circ g)^{n_{k}}(z_{0})|\rightarrow\infty.$
So, $|(f\circ g)^{m_{k}}(z_{0})|=|f^{m_{k}}\circ g^{m_{k}}(z_{0})|<R$ for some $R.$ As $f$ does not have any finite asymptotic value, so $|g^{m_{k}}(z_{0})|<R_{1}$ for some $R_{1}$. Similarly for $f,$ since they are commuting so, $|f^{m_{k}}(z_{0})|<R_{2}$ for some $R_{2}$.
Now, $|(f\circ g)^{n_{k}}(z_{0})|=|f^{n_{k}}\circ g^{n_{k}}(z_{0})|\rightarrow\infty.$ We have 3 possible cases: $|g^{n_{k}}(z_{0})|\rightarrow\infty$ or $|g^{n_{k}}(z_{0})|\rightarrow R_{3}$ for some $R_{3}$ or $g^{n_{k}}(z_{0})$ exhibits bungee behaviour. If we have the second case, it means $z_{0}\in K(g)$ (This will imply all subsequences of ($g^{j}(z_{0})$) is bounded where $j$ is an arbitrary index). To understand this, we take $w_{0}\in K(g)$. We observe $g^{j}\circ f^{2j}\circ g^{j}(w_{0})\in g^{j}\circ f^{2j}(K(g))\subset g^{j}(K(g))$ which is bounded for $j\rightarrow\infty$. Hence, this case isn’t possible. Now, when we have the third and the first case, then both essentially states that $z_{0}\in BU(g)$. From (2.2), we have $g^{n_{1k}}\circ f^{2n_{1k}}\circ g^{n_{1k}}(z_{0})$ exhibits bungee behaviour which is as expected. And due to commutation, we also conclude that $z_{0}\in BU(f)$ as well. Hence, $BU(f\circ g)\subset BU(f)\cap BU(g)$.
Given the preceding proof, we can ask a question:
Question: What can we say about points of the set $BU(f)\cup BU(g)\setminus BU(f)\cap BU(g),$ where will it be situated?
Answer: As the functions $f$ and $g$ do not have any finite asymptotic values, therefore, points of the set $BU(f)\cup BU(g)\setminus BU(f)\cap BU(g)$ cannot intersect with $K(f\circ g).$ Also, $I(f\circ g)\subset I(f)\cup I(g)$ [14] and as $(BU(f)\cup BU(g))\cap(I(f)\cup I(g))=\emptyset,$ it means that $BU(f)\cup BU(g)\setminus BU(f)\cap BU(g)$ cannot intersect with $I(f\circ g)$ either. This implies that the only place it will be contained is $BU(f\circ g)$. However, if it happens, then such points must be contained in $BU(f)\cap BU(g)$ implying that no such points exist. This is only possible if $BU(f)=BU(g)$.
Similar question can be asked for filled-in Julia sets as well, for, it was shown in Lemma 2.3 that $K(f\circ g)\subset K(f)\cap K(g)$. And the answer is: Let $w_{0}\in K(f\circ g)\ \backslash\ K(f)\cap K(g)$. These points won’t intersect with $I(f\circ g)$ and neither with $BU(f\circ g)$ as $BU(f\circ g)=BU(f)=BU(g)$ and $BU(f\circ g)\cap K(f)\cup K(g)=\emptyset.$ Similarly, for intersection of filled-in Julia sets with escaping sets. Hence, we conclude no such points are possible. Therefore, $K(f\circ g)=K(f)=K(g)$ when there are no finite asymptotic values for these commuting functions.
So given the following results:
$K(f\circ g)$ $=K(f)=K(g)$
and
(2.1) $BU(f\circ g)$ $=BU(f)=BU(g)$
one can conclude that $I(f\circ g)=I(f)=I(g)$, for we know that $K(f)\cup I(f)\cup BU(f)=\mathbb{C}=K(f\circ g)\cup I(f\circ g)\cup BU(f\circ g)=K(g)\cup I(g)\cup BU(g)$ and from the previous results (2) we can conclude that $I(f\circ g)=I(f)=I(g)$. ∎
The following examples illustrate the results established above.
Example 2.5.
Consider $f(z)=1+z+e^{-z}$(Fatou map) and $g(z)=1+z+e^{-z}+2\pi i.$ It can be easily checked that $f$ and $g$ do not have any finite asymptotic values
One may also consider the following functions for illustration of above result:
$f_{1}(z)=1+z+e^{z}$ which would be commuting with $g_{1}(z)=1+z+e^{z}+2\pi i$; $f_{2}(z)=z+sin(z)$ which would be commuting with $g_{2}(z)=z+sin(z)+2\pi.$
3. Some results for functions of class $\mathcal{B}$
Now, we will prove some results when these functions indeed have some finite asymptotic values, and to simplify, we will focus on the functions of the Eremenko-Lyubich class $\mathcal{B}$. These functions do not have a Baker domain; therefore, the escaping set does not intersect with the Fatou set. In addition, they do not have a finite intersection of the Bungee set with the wandering domain because the wandering domain of such functions does not have (sub)sequences that tend to $\infty$.
Therefore, the Julia set for functions of class $\mathcal{B}$ comprises of $RP(f),BU(f)$ and $I(f)$; that is, $J(f)=BU(f)\cup I(f)\cup RP(f)$. For commuting functions $f$ and $g$, we have $J(f\circ g)\supset J(f)\cup J(g)$ [10, 11]. However, when these are commuting functions of class $\mathcal{B}$ we have, $J(f)=J(g)=J(f\circ g)$ [10]. We also know that $g(J(f))=J(f)$ and $\overline{RP(f)}=J(f),\overline{I(f)}=J(f)$ [4, 6].
Proposition 3.1.
The bungee set for functions of class $\mathcal{B}$ does not intersect with Fatou set.
Proof.
[…]
This also means that $BU(f)\subset J(f)$ only for functions of class $\mathcal{B}.$ Also, it is easy to see that $BU(f)$ is completely invariant by the complete invariance of $I(f)$ and $K(f)$. As a result, by the minimality property of Julia set, $\overline{BU(f)}=J(f)$.
The escaping set of a function is in general, neither open, nor closed [5]. We have similar kind of result for the bungee set.
Proposition 3.2.
The bungee set for functions of class $\mathcal{B}$ is neither open nor closed.
Proof.
We can easily prove the complete invariance of bungee set by noting that $BU(f)=\mathbb{C}\backslash(I(f)\cup K(f))$. Now, if we assume that bungee set is closed then it will be a closed completely invariant set. This implies that $J(f)\subset BU(f)$ (because $J(f)$ is the smallest closed completely invariant set) which is a contradiction. Furthermore, if bungee set is open, it will imply that it has interior to it which is also not possible (because $BU(f)\subset J(f)$ and having interior of bungee set means Julia set also has interior which is not possible for an entire function $f$ for which $F(f)\neq\emptyset$ [3]). Hence, bungee set is neither open nor closed. ∎
Theorem 3.3.
Functions of class $\mathcal{B}$ which are periodic translate of each other also have equal bungee sets, escaping sets and filled-in Julia sets.
Proof.
We take $f(z)$ and $g(z)=f(z)+C$ where $C$ is the periodic translate of $f$ (more generally, we can take $g=f^{l}(z)+C$, for some $l\in\mathbb{N}$). Then $g^{n}(z)=f^{n}(z)+C$ which is almost commuting. Here we can see that if a point is bounded under $f^{n},$ it is also bounded under $g^{n}.$ Similarly, if a point is escaping under $f^{n},$ it is also escaping under $g^{n}$. Hence, $BU(f)=BU(g),I(f)=I(g),K(f)=K(g)$. ∎
Such functions can be viewed as almost commuting type with a constant shift.
Now moving forward, we will study functions that are exactly commuting of class $\mathcal{B}$ and having finite asymptotic values.
Lemma 3.4.
$g(BU(f))\subset BU(f)\cup RP(f)$ for functions in class $\mathcal{B}.$
Proof.
If we recall the definition of a bungee point, there must exist a subsequence which is bounded and another subsequence which is unbounded. Now it is obvious that when the subsequence which is bounded $f^{m}(z_{0})\rightarrow l_{0}$ (say) and when acted upon by $g$ will also keep it bounded (since $f,g$ are entire functions). Let us assume that there is only one such bounded subsequence and $l$ unbounded subsequence which are as mentioned below.
$f^{m}(z_{0})\rightarrow l_{0}$
$f^{n_{1}}(z_{0})\rightarrow\infty$
$f^{n_{2}}(z_{0})\rightarrow\infty$
$\vdots$
$f^{n_{l}}(z_{0})\rightarrow\infty$
Now, there are two possible cases when $g$ acts on those unbounded subsequences.
Case 1. All these subsequence gets bounded, i.e.,
$g(f^{n_{1}}(z_{0}))\rightarrow w_{1}\text{ which also means }f^{n_{1}}(g(z_{0}))\rightarrow w_{1}$
$g(f^{n_{2}}(z_{0}))\rightarrow w_{2}\text{ which also means }f^{n_{2}}(g(z_{0}))\rightarrow w_{2}$
$\vdots$
$g(f^{n_{l}}(z_{0}))\rightarrow w_{l}\text{ which also means }f^{n_{l}}(g(z_{0}))\rightarrow{w_{l}}$
If this happens, then $g(z_{0})\in$ $K(f)=F(f)\cup RP(f)$. As $z_{0}\in J(f)$ as well and we know that $g(J(f))=J(f)$ hence, $g(z_{0})\in RP(f)$ for the above case. Now, let us understand what does this mean by $g(z_{0})\in RP(f)$: It means that $f^{p}(g(z_{0}))=g(z_{0})$ for some $p\in\mathbb{N}.$ This also means $g(f^{p}(z_{0}))=g(z_{0})$. Because $f$ and $g$ are not invertible functions we have $f^{p}(z_{0})\not=g(z_{0})$. So, $g(z_{0})\in RP(f)\ \&\ z_{0}\in BU(f)$ can be possible.
Case 2. When at least one of the subsequence still remains unbounded. Let that be the $k-$th subsequence i.e, $g(f^{n_{k}}(z_{0}))\rightarrow\infty$ i.e. $f^{n_{k}}(g(z_{0}))\rightarrow\infty$. $g(f^{m}(z_{0}))\rightarrow l_{1},$ i.e. $f^{m}(g(z_{0}))\rightarrow l_{1}$. This shows that there are at least two subsequences and in this case $g(z_{0})$ belongs to $BU(f)$. Therefore, $g(BU(f))$ is contained in either $BU(f)$ or $RP(f)$. ∎
Lemma 3.5.
$g(I(f))\subset I(f)\cup RP(f)$ for commuting functions of class $\mathcal{B}$
Proof.
Suppose $z_{0}\in I(f)$ and let $\lim_{n\rightarrow\infty}g(f^{n}(z_{0}))\rightarrow l_{0}$ (say) i.e., $f^{n}(g(z_{0}))\rightarrow l_{0}.$ This means $g(z_{0})\in K(f).$ But $z_{0}\not\in K(f)$ as it belongs to $I(f)$. So $g(z_{0})\in RP(f)$ for the above case. Now, suppose $\lim_{n\rightarrow\infty}g(f^{n}(z_{0}))\rightarrow\infty$ i.e., $f^{n}(g(z_{0}))\rightarrow\infty$. Hence, $g$ sends $I(f)$ to either $RP(f)$ or $I(f)$. So, $g(I(f))\subset I(f)\cup RP(f)$. ∎
Proposition 3.6.
$BU(f\circ g)\subset BU(f)\cup BU(g)$ for commuting functions of class $\mathcal{B}$.
Proof.
Let $z_{0}\in BU(f\circ g)\ \backslash(BU(f)\cup BU(g))$. This implies that $z_{0}\in K(f)\cup K(g)$ or $z_{0}\in I(f)\cup I(g)$. By definition of Bungee set, there are subsequences $\{n_{k}\}$ and $\{m_{k}\}$ tending to $\infty$ satisfying
$|f^{n_{k}}(g^{n_{k}}(z_{0}))|\rightarrow\infty$
$|f^{m_{k}}(g^{m_{k}}(z_{0}))|\rightarrow l_{0}\ \ \text{for some }l_{0}.$
First, we assume that $z_{0}\in RP(g)$ (there is a clear contradiction in the $F(g)$ case because $F(g)=F(f\circ g)$), and as a result, $f^{n}(RP(g))\subset RP(g)$ and $g^{n}(f^{n}(RP(g)))$ which will be bounded. As can be seen, there are no subsequences that are unbounded. Similarly, there are no such points in $K(f)$. Therefore, $z_{0}\not\in K(f)\cup K(g)$.
Now, we assume $z_{0}\in I(g)$ which means there are two possible cases
(3.1) $|f(g^{n})(z_{0})|\rightarrow l\ \text{for some}\ l$
(3.2) $|f(g^{n})(z_{0})|\rightarrow\infty$
where $n$ denotes an arbitrary index unrelated to $n_{k}$ and $m_{k}$. If we have (3.2), then from the array of sequences argument, we see that all subsequences tend to infinity. If we have (3.1), there are further possible cases: $l\in I(f)$ or $l\in K(f)$ or $l\in BU(f)$. Again, if we have the first two cases, then we observe that entire composite sequence $(f\circ g)^{n}$ is (un)bounded. The third case is also not possible. This is because using the previous Lemma (3.4), we have $|f^{n_{k}}(g^{n_{k}})(z_{0})|$ or $|f^{m_{k}}(g^{m_{k}})(z_{0})|$ which equals $|g^{n_{k}}(f^{n_{k}})(z_{0})|$ or $|g^{m_{k}}(f^{m_{k}})(z_{0})|$ will be either completely bounded or unbounded. Therefore, $z_{0}\not\in I(g)$. On similar lines, $z_{0}\not\in I(f)$. Therefore, $z_{0}\not\in I(f)\cup I(g)$. Thus, we see that there are no points s.t, $z_{0}\in BU(f\circ g)\ \backslash(BU(f)\cup BU(g))$. ∎
Corollary 3.7.
For commuting functions of class $\mathcal{B},(I(f)\cup I(g))\cap BU(f\circ g)=\emptyset$
Proof.
We can follow the similar lines of proof that is done in the preceeding proposition (3.6). Let $z_{0}\in I(f)$ (say). Then $g(f^{n}(z_{0}))\rightarrow\infty$ or $L$. If $g(f^{n}(z_{0}))\rightarrow\infty,$ then by array of sequences argument, we conclude that $g^{n}\circ f^{n}(z_{0})\not\rightarrow M$ for some $M$ which means that we do not have a subsequence that remains bounded upon iteration of $f\circ g$. Moreover, if $g(f^{n}(z_{0}))\rightarrow L$ then $g(z_{0})\in RP(f)$ and therefore, $f^{n}(g^{n}(f^{n}(z_{0})))\rightarrow L_{1}$ (say). Thus, we never get a subsequence that tends to $\infty$.∞ . Similarly, it can be argued when $z_{0}\in I(g)$. ∎
Revision 3.8.
Revised proof of $I(f\circ g)\subset I(f)\cup I(g)$ for commuting functions of $\mathcal{B}.$ (considering the existence of Bungee set)
Proof.
The motivation for the proof comes from [14]. Let $z_{0}\in I(f\circ g)\ \backslash(I(f)\cup I(g))$ i.e., $z_{0}\in RP(f)\cup RP(g)$ or $z_{0}\in BU(f)\cup BU(g)$. By definition of escaping set: $f^{n}\circ g^{n}(z_{0})=(f\circ g)^{n}(z_{0})\rightarrow\infty$. Assume $z_{0}\in BU(f)$, then $g(BU(f))$ goes to either $BU(f)$ or $RP(f)$ (3.5). When it goes to $RP(f)$, we know $g(RP(f))$ will remain in $RP(f)$ and upon further iteration of $f\circ g$ will stay bounded i.e $(g\circ f)^{n}(z_{0})\rightarrow L$ as $n\rightarrow\infty$ which is a contradiction. If it goes to $BU(f)$ then again, its the same case as discussed before when $g$ acts on it. In the end, it will either be an entirely bounded sequence or there would be (un)bounded subsequence. The later case arises only when $g^{n}(BU(f))\subset BU(f)\ \forall n\in\mathbb{N}$ (where $n$ is some arbitrary index). But by definition escaping sets must have all the subsequences tending to infinity. Hence, we arrive at a contradiction. On similar lines, it can be shown when $z_{0}\in BU(g)$ we again arrive at a contradiction. So, $z_{0}\not\in BU(f)\cup BU(g)$.
Now the case where $z_{0}\in RP(f)\cup RP(g)$. Assuming $z_{0}\in RP(f)$ and as discussed earlier we will have all bounded subsequences of $f^{n}\circ g^{n}(z_{0})$ which is a contradiction. Similar arguments can be given when $z_{0}\in RP(g).$ Therefore $z_{0}\not\in RP(f)\cup RP(g).$ Hence, $z_{0}\in I(f\circ g)\ \backslash(I(f)\cup I(g))$ is not possible. Thus, $I(f\circ g)\subset I(f)\cup I(g)$. ∎
Proposition 3.9.
$RP(f\circ g)\supset RP(f)\cup RP(g)$ for commuting functions of class $\mathcal{B}$.
Proof.
Let us consider a point $z_{0}\in RP(f)$. Then, $f^{n}(z_{0})$ is bounded and contained in $RP(f)$. Also, $g^{n}(RP(f))\subset RP(f)$, therefore $f^{n}(g^{n}(f^{n}(z_{0})))$ is also bounded as $n\rightarrow\infty.$ (Observe that, $z_{0}$ can’t be in $F(f\circ g)$ as $F(f\circ g)=F(f)$). Hence, $z_{0}\in RP(f\circ g)$ as well. On similar lines, if $w_{0}\in RP(g)$, then $w_{0}\in RP(f\circ g)$.
Now the question is can there be points not in $RP(f)\cup RP(g)$ but in $RP(f\circ g)$? To answer that let us go through (3.4) problem once again. When suppose $g(BU(f))$ goes to $RP(f)$ and upon further iteration with $(g\circ f)^{n}(z_{0})$, this will be bounded means $z_{0}\in RP(f\circ g)$. Similarly, there will be some points from $I(f)$ that will be bounded upon $g^{n}\circ f^{n}(z_{0})$ as $n\rightarrow\infty$ as discussed in (3.5). Again it has to be in $BU(g)$ and $I(g)$. So, this implies $RP(f\circ g)\supset RP(f)\cup RP(g)$. ∎
There’s a simpler proof for this which goes as follows:
$BU(f\circ g)\cup I(f\circ g)\subset BU(f)\cup BU(g)\cup I(f)\cup I(g)$
Taking complement on both sides
$BU(f\circ g)^{c}\cap I(f\circ g)^{c}\supset BU(f)^{c}\cap BU(g)^{c}\cap I(f)^{c}\cap I(g)^{c}$
(3.3) $\implies RP(f\circ g)\supset RP(f)\cup RP(g)$
Remark 3.10.
For $f\in\mathcal{B},$ we have $J(f)=\overline{I(f)}=\overline{BU(f)}=\overline{RP(f)}=I(f)\cup BU(f)\cup RP(f).$
4. Illustrations:
We now illustrate $BU(f\circ g)\subset BU(f)\cup BU(g)$ with an example.
Consider the functions $f(z)=z^{2}$ and $g=1/z^{2}$. Observe that $f\circ g(z)=1/z^{4}$. $BU(f)=\emptyset$, $BU(g)=\{z:|z|<1\cup|z|>1\}$ and $BU(f\circ g)=\{z:|z|<1\cup|z|>1\}.$ Clearly, $BU(f\circ g)\subset BU(f)\cup BU(g)$.
Now, we discuss exponential functions which are not completely exponential type. Some examples are (see [15]): let us take the case where $f=p(z)\cdot e^{g(z)}$ where $p(z),\ g(z)$ are polynomials and $g(z)=a\cdot f(z)$ for some $a\in\mathbb{C}$. $f(z)=z\cdot e^{a^{(k+1)}\cdot z^{k}},\ k\in\mathbb{N}$ $g(z)=a\cdot z\cdot e^{a^{(k+1)}\cdot z^{k}}$.
$f\circ g(z)=f(g(z))=g(z)\cdot e^{a^{(k+1)}\cdot g^{k}(z)}=aze^{a^{(k+1)}z^{k}}e^{a^{(2k+1)}z^{k}e^{ka^{(k+1)}z^{k}}}$
$g\circ f(z)=g(f(z))=a\cdot f(z)\cdot e^{a^{k}\cdot f^{k}(z)}=aze^{a^{k}z^{k}}e^{a^{2k}z^{k}e^{ka^{k}z^{k}}}.$ Now, for $f\circ g=g\circ f,$ we need to have $a^{2k+1}=a^{k}\implies a^{k}(a^{k+1}-1)=0$ which means $a$ is $(k+1)$-root of unity except ‘1’ itself because that’s trivial.
Now, for the simplest case let us take: $k+1=2\implies a=\pm 1$ Since a=1 is a trivial case so let us take $a=-1$, $f(z)=ze^{-z^{2}};\ g(z)=-ze^{-z^{2}}$. Here as the orbit of any point under either of the function tends to $\infty,f(z)\rightarrow 0,g(z)\rightarrow 0$ and $z=0$ is a fixed point of $f(z),\ g(z)$ since $f(0)=0$ and $g(0)=0$. Now for the multiplier $\lambda$, $f^{\prime}(z)=e^{-z^{2}}(1-2z^{2}),f^{\prime}(0)=1$ which is an indifferent fixed point. And $11^{1}=$ 1 means $0\in J(f)$ or $J(g)$ (since both of them are equal). As $\infty$ is a special case and belongs to $J(f)$ so, this proves that a sequence which tends to $\infty$ if it asymptotes to a finite asymptotic value which is a member of the Julia set of $f$. Moreover, the function which are of the form $f=p(z)e^{g(z)}$ which asymptotes because it has exponential term, means the asymptotic value is ${}^{\prime}0^{\prime}$ and it is in $J(f)$ (rationally-indifferent fixed point).
Now some illustrations for the Bungee set:
Let $f(z)=ze^{z^{2}};\ g(z)=-ze^{z^{2}}$ $\implies f^{2}=ze^{z^{2}}e^{z^{2}e^{2z^{2}}}\ \&\ g^{2}=ze^{z^{2}}e^{z^{2}e^{2z^{2}}}.$ $f\circ g=-ze^{z^{2}}e^{z^{2}e^{2z^{2}}}=g\circ f$. And $(f\circ g)^{2}=g^{4}=f^{4}$(Because as $(f\circ g)^{2}$ is done it does away with the negative term and rest terms are exactly same). So, $BU(f^{4})=BU(g^{4})=BU((f\circ g)^{2})\implies BU(f)=BU(g)=BU(f\circ g)$ (NOTE: If we have $f(z)=ze^{-z^{2}};\ g(z)=-ze^{-z^{2}}$, they also commute). We may also take $f(z)=ze^{a^{(k+1)}z^{k}}$, $g(z)=aze^{a^{(k+1)}z^{k}}$ and $a^{(k+1)}=1$ for that matter). And this complies with the union relation with the composition.
Now, for the general case: Since $a$ is $k$-th root of unity, means $a^{k}=1$. So, $f(z)=ze^{a^{k}z^{k}}=ze^{z^{k}}$ and $g(z)=aze^{z^{k}}$. Now $g^{k}(z)=a^{k}\cdot f^{k}(z)=f^{k}(z)\implies BU(f)=BU(g)$. Also, $(f\circ g)^{k}(z)=a^{k}f^{2k}(z)=f^{2k}(z)$. So, $BU(f\circ g)=BU(f)=BU(g)$. Analogously, if we take $f(z)=ze^{a^{(k+1)}z^{k}}$ and $\ g(z)=af(z)=aze^{a^{(k+1)}z^{k}}$, where again $a^{k+1}=1$, we can simplify this further as $f(z)=ze^{az^{k}}$ and $g(z)=aze^{az^{k}}$, now $g^{2}(z)=af(af(z)))=a^{2}f^{2}(z)$. So, like this $g^{k+1}(z)=a^{(k+1)}f^{k+1}(z)$. Also $f\circ g(z)=af^{2}(z)=g\circ f(z)$. Now, if we do $(f\circ g)^{2}(z)=a^{2}f^{4}(z)$. So, $(f\circ g)^{k+1}(z)=a^{(k+1)}f^{2(k+1)}(z)=f^{2(k+1)}(z)$. Here also we can see that $BU(f\circ g)=BU(f)=BU(g)$.
Thus, functions which are of the form $f(z)=ze^{a^{k}z^{k}}$ where $a$ is $k$-th root of unity or $f(z)=ze^{a^{k+1}z^{k}}$, where $a$ is $(k+1)$-th root of unity and $g=af$ have the property that their bungee sets are same. Moreover, their escaping set and repelling periodic point sets are also the same. $J(f)=J(g)$ also holds true because we are taking $f=p(z)e^{g(z)}$ as discussed in [15].
5. Conclusion:
We have shown that entire functions with no finite asymptotic values which are also commuting i.e, $f\circ g=g\circ f$ satisfies:
$K(f)=K(g),\ I(f)=I(g),\ BU(f)=BU(g)$ and therefore $K(f\circ g)=K(f),\ BU(f\circ g)=BU(f),\ I(f\circ g)=I(f)$.
And for the functions of class $\mathcal{B}$ which are commuting
$RP(f\circ g)\supset RP(f)\cup RP(g)$;
$BU(f\circ g)\subset BU(f)\cup BU(g)$;
$I(f\circ g)\subset I(f)\cup I(g)$.
There are some other special cases where we have equality among them like when $f$ is an exponential family and cases like $g=f+C$ where $C\in\mathbb{C}$ which are periodic translate of $f$. Also, when functions are from class $\mathcal{B}$, $(BU(f)\cup BU(g))\cap RP(f\circ g)\neq\emptyset$ because $g(BU(f))\subset RP(f)\cup BU(f)$. Similarly, $(I(f)\cup I(g))\cap RP(f\circ g)\neq\emptyset$ because $g(I(f))\subset RP(f)\cup I(f)$. Also, there are no points which are in $I(f)\cup I(g)$ as well as in $BU(f\circ g)$.

### v3

@@
 1. Introduction
-When we iterate a homomorphic function $f:\mathbb{C}\rightarrow\mathbb{C}$, $n-$times, we observe that the points of the complex plane move differently, which we call the dynamics of the point where we denote $f^{n}$ to be $n-$th iteration of a function, where $n\in\mathbb{N}$.
+When we iterate a holomorphic function $f:\mathbb{C}\rightarrow\mathbb{C}$, $n-$times, we observe that the points of the complex plane move differently, which we call the dynamics of the point where we denote $f^{n}$ to be $n-$th iteration of a function, where $n\in\mathbb{N}$.
 As, not all points have same dynamics, we therefore have different classifications of points of the complex plane $\mathbb{C}$; one classification is based on the nature of neighbourhood of a point $z_{0}$ (say); they are said to be in the Fatou set $F(f)$ (set of points whose neighbourhood exhibits stable behaviour) and Julia set $J(f)$ (set of points whose neighbourhood exhibits chaotic behaviour).
 This was broadly the classification we generally see when we talk about dynamics of points in complex plane.
-There is yet another classification of these points based on the nature of their orbits; viz., the Filled in Julia set $K(f)$ (set of points whose orbits are bounded), escaping set $I(f)$ (set of points whose orbits tend to infinity), and bungee set $BU(f)$ (the orbits of such points neither remain bounded nor tend to infinity; only there exists a subsequence that remains bounded and another one that is unbounded).
+There is yet another classification of these points based on the nature of their orbits; viz., the filled- in Julia set $K(f)$ (set of points whose orbits are bounded), escaping set $I(f)$ (set of points whose orbits tend to infinity), and bungee set $BU(f)$ (the orbits of such points neither remain bounded nor tend to infinity; only there exists a subsequence that remains bounded and another one that is unbounded).
 These sets are disjoint among themselves and their union $BU(f)\cup I(f)\cup K(f)=\mathbb{C}$.
@@
 Knowing all these sets for an entire function, we were curious about what happens when we compose two functions and how these sets relate to the composition with those of individual functions.
-We tried to see what happens to the dynamics of transcendental functions under compositions as composition of two transcendental functions is also transcendental.
+We tried to see what happens to the dynamics of transcendental functions under compositions as composition of two transcendental functions is again transcendental.
 Simplifying the situation further, we have taken transcendental functions $f,g$ of Eremenko Lyubich class $\mathcal{B}$ which means that the functions of this class have a bounded set of singular values (i.e., the set of all asymptotic values and critical values and their finite limit points is bounded).
@@
 The bungee set of composed functions is contained in the union of the bungee set of functions used for composition.
-The set of repelling periodic and indifferent fixed points of the composed function $RP(f\circ g)$ contains the union of repelling periodic and indifferent fixed points of these functions; that is, $RP(f\circ g)\supset RP(f)\cup RP(g)$.
+The set of repelling periodic and indifferent fixed points of the composed function $RP(f\circ g)$ contains the intersection of repelling periodic and indifferent fixed points of these functions; that is, $RP(f\circ g)\supset RP(f)\cap RP(g)$.
 We know some results for exponential functions which belong to Speiser class $\mathcal{S}$ (such class of functions has finite number of elements in their singular set) [18].
 Also, $\mathcal{S}\subset\mathcal{B}$.
-$f=e^{\lambda z},\lambda\neq 0$, a function $g$ that commutes under composition arewith $f$ is some finite iteration of $f$ itself, that is, $g=f^{n}$ for some $n\in\mathbb{N}$ [2]; therefore, for such a case, all sets of composed functions are equal to the functions taken for composition.
+$f=e^{\lambda z},\lambda\neq 0$, a function $g$ that commutes under composition with $f$ is some finite iteration of $f$ itself, that is, $g=f^{n}$ for some $n\in\mathbb{N}$ [2]; therefore, in this case, the result clearly holds.
 1.1. Observations
@@
   3. (3)
-    Both sets of composite function is contained in union of individual function (to be proved in this paper).
+    Both sets of composite function is contained in union of individual functions (to be proved later on).
 2. Main results
-To begin with the commuting entire functions, we first understand the nature of these sets when functions have no asymptotic values in their singular set.
+To begin with the commuting entire functions, we first understand the nature of these sets when functions have no finite asymptotic values in their singular set.
 Theorem 2.1.
@@
 […]
-But assuming $g(z)$ does not have any finite asymptotic value, so $|g(f^{n}(z_{0})|\rightarrow\infty$.
-Since $f\,\&\,g$ are commuting so, $|g(f^{n}(z_{0})|$ = $|f^{n}(g(z_{0})|\rightarrow\infty.$ This implies that $g(z)$ carries $I(f)$ inside itself or in other words, $I(f)$ is forward invariant under $g$.
-It is easy to see that $g^{-1}(I(f))\subset I(f)$.
-Hence, $I(f)$ is completely invariant under $g$.
-As $f\,\&\,g$ are commuting, their roles can be interchanged and hence, $I(g)$ is completely invariant under $f$ as well.
-Hence, $I(f)$ and $I(g)$ is completely invariant under $f\circ g$ as well.
+Theorem 2.2.
+Suppose $f$ and $g$ are commuting entire functions having no finite asymptotic values. Then $K(f)$ and $BU(f)$ are completely invariant under $g$ and vice versa.
+Proof.
+$g(K(f))$ is, in general forward invariant, because if we assume a point $w\in K(f)$, then we have $|f^{n}(w)|<R$ for some $R.$ It is clear that $|g(f^{n}(w))|=|f^{n}(g(w))|<R_{1}$ for some $R_{1}$ which implies that $g(w)\in K(f)$, that proves the result.
+Now, let $z_{0}\in g^{-1}(K(f)).$ This implies that $g(z_{0})\in K(f)$ and $|f^{n}(g(z_{0}))|<L$ for some $L$ and because of commutation, $|g(f^{n}(z_{0}))|<L.$ As these functions do not have any finite asymptotic values, $|f^{n}(z_{0})|<L_{1}$ for some $L_{1}$, which means $z_{0}\in K(f)$.
+Thus, we have that $K(f)$ is completely invariant under $g$ when the functions have no finite asymptotic values.
+Similarly $K(g)$ is also completely invariant under $f$ as well.
+Now bungee set, $BU(f)=\mathbb{C}\backslash(I(f))\cup K(f)).$ We know from preceding argument $K(f)$ is completely invariant under $g$ and from Theorem 2.1 $I(f)$ is completely invariant under $g.$ Hence, $BU(f)$ is also completely invariant under $g$.
+Similarly, $BU(g)$ will be completely invariant under $f$.
 ∎
-Theorem 2.2.
-Suppose $f$ and $g$ are commuting entire functions having no asymptotic values. Then $K(f)$ and $BU(f)$ are completely invariant under $g$.
-Proof.
-$g(K(f))$ is, in general forward invariant, because if we assume a point $w\in K(f)$, then we have $|f^{n}(w)|<R$ for some $R.$ It is clear that $|f^{m_{k}}(g(w))|<R_{1}$ for some $R_{1}$ which implies that $g(w)\in K(f)$, thereby proving the result.
-Now, let $z_{0}\in g^{-1}(K(f)).$ This implis that $g(z_{0})\in K(f)$ and $|f^{n}(g(z_{0}))|<L$ for some $L$ and because of commutation, $|g(f^{n}(z_{0}))|<L.$ As these functions do not have any finite asymptotic values, $|f^{n}(z_{0})|<L_{1}$ for some $L_{1}$, which means $z_{0}\in K(f)$.
-Thus, we have that $K(f)$ is completely invariant under $g$ when the functions have no finite asymptotic values.
-Now bungee set, $BU(f)=\mathbb{C}\backslash(I(f))\cup K(f)).$ We know that $I(f)$ and $K(f)$ are completely invariant under $g$ (2.1) and hence, $BU(f)$ is also completely invariant under $g$. ∎
 Lemma 2.3.
-The filled-in Julia set of the composite function $f\circ g$ cannot be larger than intersection of filled in Julia set of the individual functions having no finite asymptotic values.
+The filled-in Julia set of the composite function $f\circ g$ cannot be larger than the intersection of filled-in Julia set of the individual functions which are commuting and having no finite asymptotic values.
 Proof.
@@
 Thus, there might be points in $K(f\circ g)$ that have a finite intersection with $I(f)\cup I(g)\cup BU(f)\cup BU(g)\cup A\cup B$ where $A:K(f)\backslash K(f)\cap K(g)$ and $B:K(g)\backslash K(f)\cap K(g)$.
-If $z_{0}\in I(f)\cup I(g)\cup BU(f)\cup BU(g)$, these points cannot be bounded as one function carries the points of other functions to itself, meaning $f^{n}\circ g^{n}(z_{0})$ as $n\rightarrow\infty$ will either remain oscillatory like bungee points nor will tend to $\infty$.
+If $z_{0}\in I(f)\cup I(g)\cup BU(f)\cup BU(g)$, then iterate of $z_{0}$ under $(f\circ g)^{n}$ cannot be bounded because $f^{n}\circ g^{n}(z_{0})$ as $n\rightarrow\infty$ will either remain oscillatory or will tend to $\infty$ (Theorem 2.2 and Theorem 2.1).
 Now, let $w_{0}\in A$ and apply $f^{m}\circ g^{2m}\circ f^{m}(w_{0})$ as $m\rightarrow\infty$, which is bounded because $g(K(f))\subset K(f)$.
-However, $g^{m}\circ f^{2m}\circ g^{m}(w_{0})$, which cannot be bounded ($f(BU(g)\subset BU(g)$ and $g^{m}\circ f^{2m}\circ g^{m}(w_{0})$ cannot be bounded.
-As these functions have no finite asymptotic value, so $g^{m}\circ f^{2m}\circ g^{m}(w_{0})$ will be unbounded in the case when $w_{0}\in I(g).$) Similarly, such arguments may be given when $w_{1}\in B$.
-Thus, we arrive at a contradiction.
-Hence, $K(f\circ g)\subset K(f)\cap K(g)$.
+Now, as $w_{0}\not\in K(g)$, and after interchanging roles of $f$ and $g$ due to commutation, we observe that $g^{m}\circ f^{2m}\circ g^{m}(w_{0})$ cannot be bounded (as $w_{0}$ will either be in $BU(g)$ or $I(g)$ and $f(BU(g)\subset BU(g)$, $f(I(g)\subset I(g)$).
+Hence, $g^{m}\circ f^{2m}\circ g^{m}(w_{0})=f^{m}\circ g^{2m}\circ f^{m}(w_{0})$ in which RHS is bounded while LHS is not which is a contradiction.
+Similar arguments may be given when $w_{1}\in B$.
+Hence, $K(f\circ g)\not\supset K(f)\cap K(g)$.
 ∎
+The next result provides condition for equality of bungee set, filled-in Julia set and escaping set for commuting functions.
 Theorem 2.4.
@@
 Proof.
-Let $z_{0}\in BU(f\circ g)$ where $f\ \&\ g$ do not have any finite asymptotic value.
-To begin with let there be two subsequences $\{m_{k}\}$ and $\{n_{k}\}$ both tending to $\infty$ such that $|(f\circ g)^{m_{k}}(z_{0})|<R$ for some $R$ and $|(f\circ g)^{n_{k}}(z_{0})|\rightarrow\infty.$
-So, $|(f\circ g)^{m_{k}}(z_{0})|=|f^{m_{k}}\circ g^{m_{k}}(z_{0})|<R$ for some $R.$ As $f$ does not have any finite asymptotic value, so $|g^{m_{k}}(z_{0})|<R_{1}$ for some $R_{1}$.
-Similarly for $f,$ since they are commuting so, $|f^{m_{k}}(z_{0})|<R_{2}$ for some $R_{2}$.
-Now, $|(f\circ g)^{n_{k}}(z_{0})|=|f^{n_{k}}\circ g^{n_{k}}(z_{0})|\rightarrow\infty.$ We have 3 possible cases: $|g^{n_{k}}(z_{0})|\rightarrow\infty$ or $|g^{n_{k}}(z_{0})|\rightarrow R_{3}$ for some $R_{3}$ or $g^{n_{k}}(z_{0})$ exhibits bungee behaviour.
-If we have the second case, it means $z_{0}\in K(g)$ (This will imply all subsequences of ($g^{j}(z_{0})$) is bounded where $j$ is an arbitrary index).
-To understand this, we take $w_{0}\in K(g)$.
-We observe $g^{j}\circ f^{2j}\circ g^{j}(w_{0})\in g^{j}\circ f^{2j}(K(g))\subset g^{j}(K(g))$ which is bounded for $j\rightarrow\infty$.
-Hence, this case isn’t possible.
-Now, when we have the third and the first case, then both essentially states that $z_{0}\in BU(g)$.
-From (2.2), we have $g^{n_{1k}}\circ f^{2n_{1k}}\circ g^{n_{1k}}(z_{0})$ exhibits bungee behaviour which is as expected.
-And due to commutation, we also conclude that $z_{0}\in BU(f)$ as well.
+We first establish that $BU(f\circ g)\subset BU(f)\cap BU(g).$ Let $z_{0}\in BU(f\circ g)$. By definition, there are two subsequences $\{m_{k}\}$ and $\{n_{k}\}$ both tending to $\infty$ s.t.,
+$|(f\circ g)^{m_{k}}(z_{0})|=|f^{m_{k}}\circ g^{m_{k}}(z_{0})|<R\ \text{for some}\ R$
+$|(f\circ g)^{n_{k}}(z_{0})|=|f^{n_{k}}\circ g^{n_{k}}(z_{0})|\rightarrow\infty$
+So, for $\{m_{k}\}$ subsequence, $|(f\circ g)^{m_{k}}(z_{0})|=|f^{m_{k}}\circ g^{m_{k}}(z_{0})|<R$ for some $R.$ As $f$ does not have any finite asymptotic value, so $|g^{m_{k}}(z_{0})|<R_{1}$ for some $R_{1}$.
+We observe same thing for $f$ as well i.e., $|f^{m_{k}}(z_{0})|<R_{2}$ for some $R_{2}$ (because $|f^{m_{k}}\circ g^{m_{k}}(z_{0})|=|g^{m_{k}}\circ f^{m_{k}}(z_{0})|<R$ as $f,g$ are commuting).
+Now for $\{n_{k}\}$ subsequence, $|(f\circ g)^{n_{k}}(z_{0})|=|f^{n_{k}}\circ g^{n_{k}}(z_{0})|\rightarrow\infty.$ We have 3 possible cases: $|g^{n_{k}}(z_{0})|\rightarrow\infty$ or $|g^{n_{k}}(z_{0})|\rightarrow R_{3}$ for some $R_{3}$ or $g^{n_{k}}(z_{0})$ itself exhibits bungee behaviour.
+If we have the second case, it means $z_{0}\in K(g)$ (as this will imply all subsequences of $g^{j}(z_{0})$ are bounded where $j$ is some arbitrary index unrelated to $m_{k}$ and $n_{k}$).
+As a result, $g^{j}\circ f^{2j}\circ g^{j}(z_{0})$ will remain bounded (because $g^{j}\circ f^{2j}(K(g))\subset g^{j}(K(g))$ which is bounded).
+Hence, this case isn’t possible and $z_{0}\not\in K(g)$.
+On similar lines, it can be seen that $z_{0}\not\in K(f)$ where similar arguments follow after interchanging roles of $f$ and $g$ due to commutation.
+Now, when we have either of the first or the third case, it means the same thing.
+Both these cases essentially implies that $z_{0}\in BU(g)$.
+From Theorem 2.2, we have $g^{n_{1k}}\circ f^{2n_{1k}}\circ g^{n_{1k}}(z_{0})$ exhibits bungee behaviour which is as expected.
+Due to commutation, we also conclude that $z_{0}\in BU(f)$ as well.
 Hence, $BU(f\circ g)\subset BU(f)\cap BU(g)$.
-Given the preceding proof, we can ask a question:
+Given the preceding proof, we ask a question:
 Question: What can we say about points of the set $BU(f)\cup BU(g)\setminus BU(f)\cap BU(g),$ where will it be situated?
@@
 This is only possible if $BU(f)=BU(g)$.
-Similar question can be asked for filled-in Julia sets as well, for, it was shown in Lemma 2.3 that $K(f\circ g)\subset K(f)\cap K(g)$.
-And the answer is: Let $w_{0}\in K(f\circ g)\ \backslash\ K(f)\cap K(g)$.
-These points won’t intersect with $I(f\circ g)$ and neither with $BU(f\circ g)$ as $BU(f\circ g)=BU(f)=BU(g)$ and $BU(f\circ g)\cap K(f)\cup K(g)=\emptyset.$ Similarly, for intersection of filled-in Julia sets with escaping sets.
+Similar question can be asked for filled-in Julia sets as well, (as it was shown in Lemma (2.3) that $K(f\circ g)\subset K(f)\cap K(g)$).
+And the answer is: Let $w_{0}\in K(f)\cup K(g)\ \backslash\ K(f)\cap K(g)$.
+These points won’t intersect with $I(f\circ g)$ (as $I(f\circ g)\subset I(f)\cup I(g)$ and $K(f)\cup K(g)\cap I(f)\cup I(g)=\emptyset$ [14]) nor with $BU(f\circ g)$ (as $BU(f\circ g)=BU(f)=BU(g)$ and $BU(f\circ g)\cap K(f)\cup K(g)=\emptyset)$ So, they will intersect with $K(f\circ g)$ which means $w_{0}\in K(f)\cap K(g)$ which is a contradiction.
 Hence, we conclude no such points are possible.
 Therefore, $K(f\circ g)=K(f)=K(g)$ when there are no finite asymptotic values for these commuting functions.
-So given the following results:
+Hence, given the following results:
 $K(f\circ g)$ $=K(f)=K(g)$
@@
 Example 2.5.
-Consider $f(z)=1+z+e^{-z}$(Fatou map) and $g(z)=1+z+e^{-z}+2\pi i.$ It can be easily checked that $f$ and $g$ do not have any finite asymptotic values
+Consider $f(z)=1+z+e^{-z}$(Fatou map) and $g(z)=1+z+e^{-z}+2\pi i.$ It can be easily checked that $f$ and $g$ do not have any finite asymptotic values.
 One may also consider the following functions for illustration of above result:
@@
 However, when these are commuting functions of class $\mathcal{B}$ we have, $J(f)=J(g)=J(f\circ g)$ [10].
-We also know that $g(J(f))=J(f)$ and $\overline{RP(f)}=J(f),\overline{I(f)}=J(f)$ [4, 6].
+We also know that $g(J(f))\subset J(f)$ and $\overline{RP(f)}=J(f),\overline{I(f)}=J(f)$ [4, 6].
 Proposition 3.1.
@@
 Lemma 3.4.
-$g(BU(f))\subset BU(f)\cup RP(f)$ for functions in class $\mathcal{B}.$
+$g(BU(f))\subset BU(f)\cup RP(f)$ for commuting functions of $\mathcal{B}.$
 Proof.
@@
 Now it is obvious that when the subsequence which is bounded $f^{m}(z_{0})\rightarrow l_{0}$ (say) and when acted upon by $g$ will also keep it bounded (since $f,g$ are entire functions).
-Let us assume that there is only one such bounded subsequence and $l$ unbounded subsequence which are as mentioned below.
+Let us assume that there is only one such bounded subsequence and $l$ unbounded subsequence which are as mentioned below:
 $f^{m}(z_{0})\rightarrow l_{0}$
@@
 Case 1. All these subsequence gets bounded, i.e.,
-$g(f^{n_{1}}(z_{0}))\rightarrow w_{1}\text{ which also means }f^{n_{1}}(g(z_{0}))\rightarrow w_{1}$
-$g(f^{n_{2}}(z_{0}))\rightarrow w_{2}\text{ which also means }f^{n_{2}}(g(z_{0}))\rightarrow w_{2}$
+$g(f^{n_{1}}(z_{0}))\rightarrow w_{1}=f^{n_{1}}(g(z_{0}))\rightarrow w_{1}$
+$g(f^{n_{2}}(z_{0}))\rightarrow w_{2}=f^{n_{2}}(g(z_{0}))\rightarrow w_{2}$
 $\vdots$
-$g(f^{n_{l}}(z_{0}))\rightarrow w_{l}\text{ which also means }f^{n_{l}}(g(z_{0}))\rightarrow{w_{l}}$
+$g(f^{n_{l}}(z_{0}))\rightarrow w_{l}=f^{n_{l}}(g(z_{0}))\rightarrow{w_{l}}$
 If this happens, then $g(z_{0})\in$ $K(f)=F(f)\cup RP(f)$.
-As $z_{0}\in J(f)$ as well and we know that $g(J(f))=J(f)$ hence, $g(z_{0})\in RP(f)$ for the above case.
+As $z_{0}\in J(f)$ and we know that $g(J(f))\subset J(f)$ hence, $g(z_{0})\in RP(f)$ for the above case.
 Now, let us understand what does this mean by $g(z_{0})\in RP(f)$: It means that $f^{p}(g(z_{0}))=g(z_{0})$ for some $p\in\mathbb{N}.$ This also means $g(f^{p}(z_{0}))=g(z_{0})$.
-Because $f$ and $g$ are not invertible functions we have $f^{p}(z_{0})\not=g(z_{0})$.
+Because $f$ and $g$ are not invertible functions we have $f^{p}(z_{0})\not=z_{0}$.
 So, $g(z_{0})\in RP(f)\ \&\ z_{0}\in BU(f)$ can be possible.
@@
 When at least one of the subsequence still remains unbounded.
-Let that be the $k-$th subsequence i.e, $g(f^{n_{k}}(z_{0}))\rightarrow\infty$ i.e.
-$f^{n_{k}}(g(z_{0}))\rightarrow\infty$.
+Let that be the $k-$th subsequence i.e., $g(f^{n_{k}}(z_{0}))\rightarrow\infty$ i.e., $f^{n_{k}}(g(z_{0}))\rightarrow\infty$.
 $g(f^{m}(z_{0}))\rightarrow l_{1},$ i.e.
@@
 Lemma 3.5.
-$g(I(f))\subset I(f)\cup RP(f)$ for commuting functions of class $\mathcal{B}$
+$g(I(f))\subset I(f)\cup RP(f)$ for commuting functions of $\mathcal{B}.$
 Proof.
@@
 $|f^{m_{k}}(g^{m_{k}}(z_{0}))|\rightarrow l_{0}\ \ \text{for some }l_{0}.$
-First, we assume that $z_{0}\in RP(g)$ (there is a clear contradiction in the $F(g)$ case because $F(g)=F(f\circ g)$), and as a result, $f^{n}(RP(g))\subset RP(g)$ and $g^{n}(f^{n}(RP(g)))$ which will be bounded.
+First, we assume that $z_{0}\in K(g)$ which in turn implies $z_{0}\in RP(g)$ (there is a clear contradiction in the $F(g)$ case because $F(g)=F(f\circ g)$), and as a result, $f^{n}(RP(g))\subset RP(g)$ and $g^{n}(f^{n}(RP(g)))$ which will be bounded.
 As can be seen, there are no subsequences that are unbounded.
@@
 Now, we assume $z_{0}\in I(g)$ which means there are two possible cases
-(3.1) $|f(g^{n})(z_{0})|\rightarrow l\ \text{for some}\ l$
-(3.2) $|f(g^{n})(z_{0})|\rightarrow\infty$
+(3.1) $|f(g^{n}(z_{0}))|\rightarrow l\ \text{for some}\ l$
+(3.2) $|f(g^{n}(z_{0}))|\rightarrow\infty$
 where $n$ denotes an arbitrary index unrelated to $n_{k}$ and $m_{k}$.
-If we have (3.2), then from the array of sequences argument, we see that all subsequences tend to infinity.
-If we have (3.1), there are further possible cases: $l\in I(f)$ or $l\in K(f)$ or $l\in BU(f)$.
+If we have case (3.2), then from the array of sequences argument, we see that all subsequences of $(f\circ g)^{n}(z_{0})$ tend to infinity.
+If we have case (3.1), there are further possible cases: $l\in I(f)$ or $l\in K(f)$ or $l\in BU(f)$.
 Again, if we have the first two cases, then we observe that entire composite sequence $(f\circ g)^{n}$ is (un)bounded.
@@
 Therefore, $z_{0}\not\in I(f)\cup I(g)$.
-Thus, we see that there are no points s.t, $z_{0}\in BU(f\circ g)\ \backslash(BU(f)\cup BU(g))$.
+Thus, we see that there are no points s.t., $z_{0}\in BU(f\circ g)\ \backslash(BU(f)\cup BU(g))$.
 ∎
@@
 By definition of escaping set: $f^{n}\circ g^{n}(z_{0})=(f\circ g)^{n}(z_{0})\rightarrow\infty$.
-Assume $z_{0}\in BU(f)$, then $g(BU(f))$ goes to either $BU(f)$ or $RP(f)$ (3.5).
-When it goes to $RP(f)$, we know $g(RP(f))$ will remain in $RP(f)$ and upon further iteration of $f\circ g$ will stay bounded i.e $(g\circ f)^{n}(z_{0})\rightarrow L$ as $n\rightarrow\infty$ which is a contradiction.
+Assume $z_{0}\in BU(f)$, then $g(BU(f))$ goes to either $BU(f)$ or $RP(f)$ using Lemma 3.4.
+When it goes to $RP(f)$, we know $g(RP(f))$ will remain in $RP(f)$ and upon further iteration of $f\circ g$ will stay bounded i.e., $(g\circ f)^{n}(z_{0})\rightarrow L$ as $n\rightarrow\infty$ which is a contradiction.
 If it goes to $BU(f)$ then again, its the same case as discussed before when $g$ acts on it.
-In the end, it will either be an entirely bounded sequence or there would be (un)bounded subsequence.
+In the end, it will either be an entirely bounded sequence or there would be an (un)bounded subsequence.
 The later case arises only when $g^{n}(BU(f))\subset BU(f)\ \forall n\in\mathbb{N}$ (where $n$ is some arbitrary index).
-But by definition escaping sets must have all the subsequences tending to infinity.
+But by definition, escaping sets must have all the subsequences tending to infinity.
 Hence, we arrive at a contradiction.
-On similar lines, it can be shown when $z_{0}\in BU(g)$ we again arrive at a contradiction.
+On similar lines, it can be shown that when $z_{0}\in BU(g)$ we again arrive at a contradiction.
 So, $z_{0}\not\in BU(f)\cup BU(g)$.
@@
 Proposition 3.9.
-$RP(f\circ g)\supset RP(f)\cup RP(g)$ for commuting functions of class $\mathcal{B}$.
+$RP(f\circ g)\supset RP(f)\cap RP(g)$ for commuting functions of class $\mathcal{B}$.
 Proof.
-Let us consider a point $z_{0}\in RP(f)$.
-Then, $f^{n}(z_{0})$ is bounded and contained in $RP(f)$.
-Also, $g^{n}(RP(f))\subset RP(f)$, therefore $f^{n}(g^{n}(f^{n}(z_{0})))$ is also bounded as $n\rightarrow\infty.$ (Observe that, $z_{0}$ can’t be in $F(f\circ g)$ as $F(f\circ g)=F(f)$).
-Hence, $z_{0}\in RP(f\circ g)$ as well.
-On similar lines, if $w_{0}\in RP(g)$, then $w_{0}\in RP(f\circ g)$.
-Now the question is can there be points not in $RP(f)\cup RP(g)$ but in $RP(f\circ g)$?
-To answer that let us go through (3.4) problem once again.
-When suppose $g(BU(f))$ goes to $RP(f)$ and upon further iteration with $(g\circ f)^{n}(z_{0})$, this will be bounded means $z_{0}\in RP(f\circ g)$.
-Similarly, there will be some points from $I(f)$ that will be bounded upon $g^{n}\circ f^{n}(z_{0})$ as $n\rightarrow\infty$ as discussed in (3.5).
-Again it has to be in $BU(g)$ and $I(g)$.
-So, this implies $RP(f\circ g)\supset RP(f)\cup RP(g)$.
-∎
-There’s a simpler proof for this which goes as follows:
 $BU(f\circ g)\cup I(f\circ g)\subset BU(f)\cup BU(g)\cup I(f)\cup I(g)$
@@
 $BU(f\circ g)^{c}\cap I(f\circ g)^{c}\supset BU(f)^{c}\cap BU(g)^{c}\cap I(f)^{c}\cap I(g)^{c}$
-(3.3) $\implies RP(f\circ g)\supset RP(f)\cup RP(g)$
+(3.3) $\implies RP(f\circ g)\supset RP(f)\cap RP(g)$
 Remark 3.10.
@@
 4. Illustrations:
-We now illustrate $BU(f\circ g)\subset BU(f)\cup BU(g)$ with an example.
+We now illustrate $BU(f\circ g)\subset BU(f)\cup BU(g)$ with some examples. Our first example is commuting rational functions.
 Consider the functions $f(z)=z^{2}$ and $g=1/z^{2}$.
@@
 $BU(f)=\emptyset$, $BU(g)=\{z:|z|<1\cup|z|>1\}$ and $BU(f\circ g)=\{z:|z|<1\cup|z|>1\}.$ Clearly, $BU(f\circ g)\subset BU(f)\cup BU(g)$.
-Now, we discuss exponential functions which are not completely exponential type.
-Some examples are (see [15]): let us take the case where $f=p(z)\cdot e^{g(z)}$ where $p(z),\ g(z)$ are polynomials and $g(z)=a\cdot f(z)$ for some $a\in\mathbb{C}$.
-$f(z)=z\cdot e^{a^{(k+1)}\cdot z^{k}},\ k\in\mathbb{N}$ $g(z)=a\cdot z\cdot e^{a^{(k+1)}\cdot z^{k}}$.
+Our next example is motivated from [15].
+Let us take the case where $f=p(z)\cdot e^{h(z)}$ where $p(z),\ h(z)$ are polynomials and $g(z)=a\cdot f(z)$ for some $a\in\mathbb{C}\setminus\{0\}$.
+For simplicity, let us take $f(z)=z\cdot e^{a^{(k+1)}\cdot z^{k}},\ k\in\mathbb{N}$ $g(z)=a\cdot z\cdot e^{a^{(k+1)}\cdot z^{k}}$.
 $f\circ g(z)=f(g(z))=g(z)\cdot e^{a^{(k+1)}\cdot g^{k}(z)}=aze^{a^{(k+1)}z^{k}}e^{a^{(2k+1)}z^{k}e^{ka^{(k+1)}z^{k}}}$
-$g\circ f(z)=g(f(z))=a\cdot f(z)\cdot e^{a^{k}\cdot f^{k}(z)}=aze^{a^{k}z^{k}}e^{a^{2k}z^{k}e^{ka^{k}z^{k}}}.$ Now, for $f\circ g=g\circ f,$ we need to have $a^{2k+1}=a^{k}\implies a^{k}(a^{k+1}-1)=0$ which means $a$ is $(k+1)$-root of unity except ‘1’ itself because that’s trivial.
+$g\circ f(z)=g(f(z))=a\cdot f(z)\cdot e^{a^{k}\cdot f^{k}(z)}=aze^{a^{k}z^{k}}e^{a^{2k}z^{k}e^{ka^{k}z^{k}}}.$ Now, for $f\circ g=g\circ f,$ we need to have $a^{2k+1}=a^{k}\implies a^{k}(a^{k+1}-1)=0,$ which means $a$ is $(k+1)$-root of unity except ‘1’ itself because that’s trivial.
 Now, for the simplest case let us take: $k+1=2\implies a=\pm 1$ Since a=1 is a trivial case so let us take $a=-1$, $f(z)=ze^{-z^{2}};\ g(z)=-ze^{-z^{2}}$.
@@
 As $\infty$ is a special case and belongs to $J(f)$ so, this proves that a sequence which tends to $\infty$ if it asymptotes to a finite asymptotic value which is a member of the Julia set of $f$.
-Moreover, the function which are of the form $f=p(z)e^{g(z)}$ which asymptotes because it has exponential term, means the asymptotic value is ${}^{\prime}0^{\prime}$ and it is in $J(f)$ (rationally-indifferent fixed point).
+Moreover, the functions which are of the form $f=p(z)e^{h(z)}$ have a finite asymptotic value ${}^{\prime}0^{\prime}$ and it is in $J(f)$ (rationally indifferent fixed point or repelling periodic point).
 Now some illustrations for the bungee set:
@@
 So, $BU(f\circ g)=BU(f)=BU(g)$.
-Analogously, if we take $f(z)=ze^{a^{(k+1)}z^{k}}$ and $\ g(z)=af(z)=aze^{a^{(k+1)}z^{k}}$, where again $a^{k+1}=1$, we can simplify this further as $f(z)=ze^{az^{k}}$ and $g(z)=aze^{az^{k}}$, now $g^{2}(z)=af(af(z)))=a^{2}f^{2}(z)$.
-So, like this $g^{k+1}(z)=a^{(k+1)}f^{k+1}(z)$.
-Also $f\circ g(z)=af^{2}(z)=g\circ f(z)$.
-Now, if we do $(f\circ g)^{2}(z)=a^{2}f^{4}(z)$.
-So, $(f\circ g)^{k+1}(z)=a^{(k+1)}f^{2(k+1)}(z)=f^{2(k+1)}(z)$.
-Here also we can see that $BU(f\circ g)=BU(f)=BU(g)$.
+Analogously, if we take $f(z)=ze^{a^{(k+1)}z^{k}}$ and $\ g(z)=af(z)=aze^{a^{(k+1)}z^{k}}$, where again $a^{k+1}=1$, we can simplify it further as $f(z)=ze^{az^{k}}$ and $g(z)=aze^{az^{k}}$, now $g^{2}(z)=af(af(z)))=a^{2}f^{2}(z)$.
+By induction, $g^{k+1}(z)=a^{(k+1)}f^{k+1}(z)$.
+Also, $f\circ g(z)=af^{2}(z)=g\circ f(z)$ and $(f\circ g)^{2}(z)=a^{2}f^{4}(z)$.
+Proceeding by induction, $(f\circ g)^{k+1}(z)=a^{(k+1)}f^{2(k+1)}(z)=f^{2(k+1)}(z)$.
+Here also, we can see that $BU(f\circ g)=BU(f)=BU(g)$.
 Thus, functions which are of the form $f(z)=ze^{a^{k}z^{k}}$ where $a$ is $k$-th root of unity or $f(z)=ze^{a^{k+1}z^{k}}$, where $a$ is $(k+1)$-th root of unity and $g=af$ have the property that their bungee sets are same.
@@
 5. Conclusion:
-We have shown that entire functions with no finite asymptotic values which are also commuting i.e, $f\circ g=g\circ f$ satisfies:
-$K(f)=K(g),\ I(f)=I(g),\ BU(f)=BU(g)$ and therefore $K(f\circ g)=K(f),\ BU(f\circ g)=BU(f),\ I(f\circ g)=I(f)$.
+We have shown that entire functions with no finite asymptotic values which are also commuting i.e., $f\circ g=g\circ f$ satisfies:
+$K(f)=K(g),\ I(f)=I(g),\ BU(f)=BU(g)$ and therefore, $K(f\circ g)=K(f),\ BU(f\circ g)=BU(f),\ I(f\circ g)=I(f)$.
 And for the functions of class $\mathcal{B}$ which are commuting
-$RP(f\circ g)\supset RP(f)\cup RP(g)$;
+$RP(f\circ g)\supset RP(f)\cap RP(g)$;
 $BU(f\circ g)\subset BU(f)\cup BU(g)$;
 $I(f\circ g)\subset I(f)\cup I(g)$.
-There are some other special cases where we have equality among them like when $f$ is an exponential family and cases like $g=f+C$ where $C\in\mathbb{C}$ which are periodic translate of $f$.
+There are some other special cases where we have equality among these sets, like when $f$ is an exponential map.
+Also, cases like $g=f+C$ where $C\in\mathbb{C}$ is the period of $f$.
 Also, when functions are from class $\mathcal{B}$, $(BU(f)\cup BU(g))\cap RP(f\circ g)\neq\emptyset$ because $g(BU(f))\subset RP(f)\cup BU(f)$.

### v4

@@
 Abstract.
-In this paper, we have discussed the dynamics of composite entire functions in terms of relationship between bungee set, repelling periodic points (to be denoted by $RP$) and rationally indifferent fixed point set.
-We have established relation between the dynamics of composite function and the functions taken for composition.
-We showed that the union of their bungee sets contains bungee set of the composite function.
-We also showed that $RP$ set of composite functions contains the $RP$ set of functions used for the composition.
-We have mostly dealt with the permutable functions of class $\mathcal{B}$.
-These results hold true when we have a composite function in which one function is a periodic translation of another.
+In this paper, we have discussed the dynamics of composite entire functions in terms of relationship between bungee set, escaping set and filled-in Julia set.
+We have established some relation between the dynamics of composition of entire functions and the functions taken for composition.
+We have shown that the union of the bungee set of two entire functions contains the bungee set of the composite function.
+In addition, it is shown that the filled-in Julia set of composite entire functions contains the filled-in Julia set of functions used for the composition.
+The results have been illustrated with several examples.
+We have mostly dealt with permutable(commuting) functions.
 Key words and phrases:
@@
 1. Introduction
-When we iterate a holomorphic function $f:\mathbb{C}\rightarrow\mathbb{C}$, $n-$times, we observe that the points of the complex plane move differently, which we call the dynamics of the point where we denote $f^{n}$ to be $n-$th iteration of a function, where $n\in\mathbb{N}$.
+When we iterate a holomorphic function $f:\mathbb{C}\rightarrow\mathbb{C}$, $n$-times, we observe that the points of the complex plane move differently, which we call the dynamics of the point.
+Here, $f^{n}$ denotes the $n$-th iteration of function $f$, where $n\in\mathbb{N}$.
 As, not all points have same dynamics, we therefore have different classifications of points of the complex plane $\mathbb{C}$; one classification is based on the nature of neighbourhood of a point $z_{0}$ (say); they are said to be in the Fatou set $F(f)$ (set of points whose neighbourhood exhibits stable behaviour) and Julia set $J(f)$ (set of points whose neighbourhood exhibits chaotic behaviour).
@@
 In [16], it was shown that if a Fatou component intersects with $BU(f)$ then that component becomes a wandering domain contained in $BU(f)$.
-$K(f)$ comprises of Fatou components in which orbits under iteration stays bounded, along with repelling periodic points and rationally indifferent periodic points that lies inside Julia set $J(f).$ It is to be noted that the bungee set and escaping set must have a finite intersection with $J(f)$ (see [5, 16]).
-The escaping set too has a finite intersection with the Fatou set when the Fatou component is a Baker domain.
-Knowing all these sets for an entire function, we were curious about what happens when we compose two functions and how these sets relate to the composition with those of individual functions.
-We tried to see what happens to the dynamics of transcendental functions under compositions as composition of two transcendental functions is again transcendental.
-Simplifying the situation further, we have taken transcendental functions $f,g$ of Eremenko Lyubich class $\mathcal{B}$ which means that the functions of this class have a bounded set of singular values (i.e., the set of all asymptotic values and critical values and their finite limit points is bounded).
-Additionally, this class does not have any intersection of the Fatou set with the bungee set, in other words there is no wandering domain which also lies inside the bungee set for $f\in\mathcal{B}.$ Since our aim was to establish a relationship between different sets of functions and their composition and it would be quite difficult to establish any relation where there is no algebraic relation among these functions, we have focused on commuting(permutable) functions, that is, $f\circ g=g\circ f$, and $g=f+C$ where $C$ is the period of $f$.
+It is to be noted that the bungee set and escaping set must have a non empty intersection with $J(f)$ (see [5, 16]).
+The escaping set too has a non empty intersection with the Fatou set when the Fatou component is a Baker domain [14], or when the Fatou set has a multiply connected wandering domain [19].
+Knowing all these sets for an entire function, we were curious about what happens when we compose two such functions and how these sets relate to the composition with those of individual functions.
+As composition of two transcendental functions is again transcendental, we have tried to analyse what happens to the dynamics of transcendental functions under compositions.
+Simplifying the situation further, we have taken transcendental functions $f,g$ in the Eremenko-Lyubich class $\mathcal{B}$ which means that the functions of this class have a bounded set of singular values (i.e., the set of all asymptotic values and critical values and their finite limit points is bounded).
+Recall the notion of an oscillatory wandering domain.
+Definition 1.1.
+Given a transcendental entire function $f$, a domain $U\subset F(f)$ is called an oscillatory wandering domain of $f$, if $U\subseteq BU(f)$, that is, $\{\infty,d\}$ is a limit function of $\{f^{n}\}$ on $U$ for some $d\in\mathbb{C}$, [17].
+Additionally, class $\mathcal{B}$ may have an intersection of the Fatou set with the bungee set, in other words there can be oscillating wandering domain which also lies inside the bungee set for $f\in\mathcal{B}.$ As our aim was to establish a relationship between different sets of functions and their compositions, it would be very difficult to establish any such relation in the absence of some algebraic relation among these functions.
+Hence, we have focused on permutable (commuting) functions, that is, functions $f,g$ satisfying $f\circ g=g\circ f.$ Also, we have considered entire functions $f$ and $g$ related by $g=f+C$ where $C$ is the period of $f$.
 The following results are proven in this paper:
 Let $f,g\in\mathcal{B}$ be commuting functions.
-The bungee set of composed functions is contained in the union of the bungee set of functions used for composition.
-The set of repelling periodic and indifferent fixed points of the composed function $RP(f\circ g)$ contains the intersection of repelling periodic and indifferent fixed points of these functions; that is, $RP(f\circ g)\supset RP(f)\cap RP(g)$.
+The bungee set of composed function is contained in the union of the bungee set of functions used for composition.
 We know some results for exponential functions which belong to Speiser class $\mathcal{S}$ (such class of functions has finite number of elements in their singular set) [20].
 Also, $\mathcal{S}\subset\mathcal{B}$.
-$f=e^{\lambda z},\lambda\neq 0$, a function $g$ that commutes under composition with $f$ is some finite iteration of $f$ itself, that is, $g=f^{n}$ for some $n\in\mathbb{N}$ [2]; therefore, in this case, the result clearly holds.
+It was shown in [2] that for $f=e^{\lambda z},\lambda\neq 0$, a function $g$ that commutes with $f$ is some finite iteration of $f$ itself, that is, $g=f^{n}$ for some $n\in\mathbb{N}$.
+Hence, in this case, the result clearly holds, in fact, we obtain $BU(f)=BU(g)=BU(f\circ g).$
 1.1. Observations
-While developing the relationships among the functions and the individual functions, we also found that the bungee set and escaping set exhibit similar characteristics for functions of class $\mathcal{B}$:
+While developing the relationships among the sets of composite functions and the individual functions, we also found that the bungee set and escaping set exhibit similar characteristics for functions of class $\mathcal{B}$:
   1. (1)
@@
 2. Main results
-To begin with the commuting entire functions, we first understand the nature of these sets when functions have no finite asymptotic values in their singular set.
+Property A : $f$ (respectively $g$) has the property that it sends orbits escaping under $g$ (respectively, $f$) to $\infty.$
+To begin with the commuting entire functions, we first understand the nature of these sets when functions satisfies Property A.
 Theorem 2.1.
-If $f$ and $g$ are commuting functions having no finite asymptotic values, then $I(f)$ is completely invariant under $g$ and vice versa.
+If $f$ and $g$ are commuting functions satisfying Property A, then $I(f)$ is completely invariant under $g$ and vice versa.
 Proof.
+Suppose $z_{0}\in I(f).$ Then, $|f^{n}(z_{0})|\rightarrow\infty$.
+Now $|g(f^{n}(z_{0})|$ can either tend to $\infty$ or be bounded.
+Using Property A, $|g(f^{n}(z_{0})|\rightarrow\infty$.
+As $f$ and $g$ are commuting so, $|g(f^{n}(z_{0}))|$ = $|f^{n}(g(z_{0}))|\rightarrow\infty.$ This implies that $I(f)$ is forward invariant under $g$.
+It is easy to see that $g^{-1}(I(f))\subset I(f)$.
+Hence, $I(f)$ is completely invariant under $g$.
+As $f$ and $g$ are commuting, their roles can be interchanged and hence, $I(g)$ is completely invariant under $f$ as well.
+∎
+Remark 2.2.
+It follows from above result that $I(f)$ and $I(g)$ are completely invariant under $f\circ g$ as well.
+Theorem 2.3.
+Suppose $f$ and $g$ are commuting entire functions satisfying Property A. Then $K(f)$ and $BU(f)$ are completely invariant under $g$ and vice versa.
+Proof.
+$K(f)$ is in general forward invariant under $g,$ because if we take a point $w\in K(f)$, then we have $|f^{n}(w)|<R$ for some $R.$ It is clear that $|g(f^{n}(w))|=|f^{n}(g(w))|<R_{1}$ for some $R_{1}$ which implies that $g(w)\in K(f)$ that proves the result.
+Now, let $z_{0}\in g^{-1}(K(f)).$ This implies that $|f^{n}(g(z_{0}))|<L$ for some $L$ and because of commutation, $|g(f^{n}(z_{0}))|<L.$ Using Property A, $|f^{n}(z_{0})|<L_{1}$ for some $L_{1}$, which means $z_{0}\in K(f)$.
+Thus, we obtain that $K(f)$ is completely invariant under $g$.
+Similarly, $K(g)$ is also completely invariant under $f$ as well.
+∎
+Theorem 2.4.
+Suppose $f$ and $g$ are commuting entire functions satisfying Property A. Then $BU(f)$ is completely invariant under $g$ and vice versa.
+Proof.
+Using the dynamical partition of the complex plane, we obtain $BU(f)=\mathbb{C}\backslash(I(f))\cup K(f)).$ We know from preceding argument that $K(f)$ is completely invariant under $g$ and from Theorem 2.1, $I(f)$ is completely invariant under $g.$ Hence, on combining these results, we conclude that $BU(f)$ is also completely invariant under $g$.
+On similar lines, $BU(g)$ will also be completely invariant under $f$.
+∎
+Remark 2.5.
+It follows from above result that $I(f),$ $I(g);K(f),K(g);$ and $BU(f),BU(g)$ are completely invariant under $f\circ g$ as well.
+The next result provides a relation between the filled-in Julia set of composite entire functions and the filled-in Julia set of the individual functions.
+Lemma 2.6.
+The filled-in Julia set of the composite function $f\circ g$ contains the intersection of the filled-in Julia set of the individual functions which are commuting and satisfying Property A.
+Proof.
+Suppose that $z_{0}\in K(f)\cap K(g)$. Consider the array of following sequences:
+$\begin{array}[]{ccccc}f(g(z_{0}))&f^{2}(g(z_{0}))&f^{3}(g(z_{0}))\cdots\\ f(g^{2}(z_{0}))&f^{2}(g^{2}(z_{0}))&f^{3}(g^{2}(z_{0}))\cdots\\ \vdots&\vdots&\vdots\\ f(g^{k}(z_{0}))&f^{2}(g^{k}(z_{0}))&f^{3}(g^{k}(z_{0}))\cdots\\$ $\vdots$$\vdots&\vdots\par\end{array}$
+Now, the first row of the array will converge to say, $l$ for, if it does not, then $g(z_{0})\in I(f)$ which implies that $z_{0}\in I(f)$, a contradiction.
+Using continuity of $g$, we can say that second row will converge to $g(l)$ and so on.
+Let $X$ be the set containing all these limit points.
+The following are the possible cases for set $X$.
+Case 1: $l$ is periodic point of $g$. In particular, suppose that $l$ is a fixed point of $g$. Hence, every sequence of the array will be bounded which shows that $z_{0}\in K(f\circ g)$.
+Case 2: $X$ is infinite and bounded. Then, by Bolzano-Weierstrass theorem, $X$ has a limit point which also shows that $z_{0}\in K(f\circ g)$.
+Case 3: $X$ is unbounded, so that there exists a subsequence which escapes to infinity. This in turn implies that $z_{0}\in I(f\circ g)\subset I(f)\cup I(g)$ (see [14]) which leads to a contradiction.
+From all the above cases, we obtain that $K(f)\cap K(g)\subset K(f\circ g).$
+The next result connects bungee set of composite entire function with that of individual functions.
+Theorem 2.7.
+Suppose $f$ and $g$ are commuting transcendental entire functions satisfying Property A. Then, $BU(f\circ g)\subset BU(f)\cap BU(g).$
+Proof.
+Suppose $z_{0}\in BU(f\circ g).$ Then, there exists two subsequences $\{n_{k}\},\{m_{k}\}$ and a constant $M>0$ such that $(f\circ g)^{n_{k}}(z_{0})\to\infty$ as $k\to\infty$ and $|(f\circ g)^{m_{k}}(z_{0})|<M\mbox{ for all }k\in\mathbb{N}$.
+It is enough to show that $z_{0}\in BU(g)$ (as the proof of $z_{0}\in BU(f)$ follows on similar lines).
+In order to show this, we prove that $z_{0}$ cannot be in $I(g)$ as well as $K(g).$ We divide the proof into several cases:
+Case(i): Suppose that $g^{n_{k}}(z_{0})\to\infty$ and $g^{m_{k}}(z_{0})\to\infty$ as $k\to\infty$.
+We can assume that $g^{n}(z_{0})\to\infty$ as $n\to\infty$ (for, if there exists a subsequence $\{l_{k}\}$ such that $g^{l_{k}}(z_{0})$ stays bounded as $k\to\infty$ then this will clearly imply that $z_{0}\in BU(g)$).
+Also, using Theorem 2.1, $f^{k}(I(g)\subset I(g)$ for every $k\in\mathbb{N}$.
+In particular, $f^{m_{k}}(z_{0})\in I(g)$ which implies that $(g\circ f)^{m_{k}}(z_{0})\to\infty$ as $k\to\infty$ which leads to a contradiction.
+Case(ii): Suppose that $g^{n_{k}}(z_{0})\to b$ and $g^{m_{k}}(z_{0})\to\infty$ as $k\to\infty$. This again shows that $z_{0}\in BU(g)$ and the result follows.
+Case(3): Suppose that $g^{n_{k}}(z_{0})\to\alpha$ and $g^{m_{k}}(z_{0})\to\beta$ as $k\to\infty.$ It follows that $z_{0}\in K(g)$ (for, if there exists a subsequence $\{l_{k}\}$ for which $g^{l_{k}}(z_{0})\to\infty$ as $k\to\infty$ then we are back to Case(2) and hence $z_{0}\in BU(g)$).
+Using Theorem 2.1, $f(K(g)\subset K(g)$.
+In particular, $f^{n_{k}}(z_{0})\in K(g)$, i.e., there exists $L>0$ such that $|g^{n}(f^{n_{k}}(z_{0}))|<L\mbox{ for all }n\in\mathbb{N}$.
+This implies that $|(f\circ g)^{n_{k}}(z_{0})|<L\mbox{ for all }k\in\mathbb{N}$ which is again a contradiction.
+Thus, from the above cases, we conclude that $z_{0}\in BU(g).$ The proof of $z_{0}\in BU(f)$ follows on similar lines.
+Hence, we obtain that $BU(f\circ g)\subset BU(f)\cap BU(g)\subset BU(f)\cup BU(g)$.
+∎
+Given the preceding proof, we ask a question:
+Question: What can we say about points of the set $K(f)\cup K(g)\setminus K(f)\cap K(g),$ where will it be situated?
 […]
-Theorem 2.2.
-Suppose $f$ and $g$ are commuting entire functions having no finite asymptotic values. Then $K(f)$ and $BU(f)$ are completely invariant under $g$ and vice versa.
-Proof.
-$g(K(f))$ is, in general forward invariant, because if we assume a point $w\in K(f)$, then we have $|f^{n}(w)|<R$ for some $R.$ It is clear that $|g(f^{n}(w))|=|f^{n}(g(w))|<R_{1}$ for some $R_{1}$ which implies that $g(w)\in K(f)$, that proves the result.
-Now, let $z_{0}\in g^{-1}(K(f)).$ This implies that $g(z_{0})\in K(f)$ and $|f^{n}(g(z_{0}))|<L$ for some $L$ and because of commutation, $|g(f^{n}(z_{0}))|<L.$ As these functions do not have any finite asymptotic values, $|f^{n}(z_{0})|<L_{1}$ for some $L_{1}$, which means $z_{0}\in K(f)$.
-Thus, we have that $K(f)$ is completely invariant under $g$ when the functions have no finite asymptotic values.
-Similarly $K(g)$ is also completely invariant under $f$ as well.
-Now bungee set, $BU(f)=\mathbb{C}\backslash(I(f))\cup K(f)).$ We know from preceding argument $K(f)$ is completely invariant under $g$ and from Theorem 2.1 $I(f)$ is completely invariant under $g.$ Hence, $BU(f)$ is also completely invariant under $g$.
-Similarly, $BU(g)$ will be completely invariant under $f$.
-∎
-Lemma 2.3.
-The filled-in Julia set of the composite function $f\circ g$ cannot be larger than the intersection of filled-in Julia set of the individual functions which are commuting and having no finite asymptotic values.
-Proof.
-We assume that $K(f\circ g)\supset K(f)\cap K(g)$.
-Thus, there might be points in $K(f\circ g)$ that have a finite intersection with $I(f)\cup I(g)\cup BU(f)\cup BU(g)\cup A\cup B$ where $A:K(f)\backslash K(f)\cap K(g)$ and $B:K(g)\backslash K(f)\cap K(g)$.
-If $z_{0}\in I(f)\cup I(g)\cup BU(f)\cup BU(g)$, then iterate of $z_{0}$ under $(f\circ g)^{n}$ cannot be bounded because $f^{n}\circ g^{n}(z_{0})$ as $n\rightarrow\infty$ will either remain oscillatory or will tend to $\infty$ (Theorem 2.2 and Theorem 2.1).
-Now, let $w_{0}\in A$ and apply $f^{m}\circ g^{2m}\circ f^{m}(w_{0})$ as $m\rightarrow\infty$, which is bounded because $g(K(f))\subset K(f)$.
-Now, as $w_{0}\not\in K(g)$, and after interchanging roles of $f$ and $g$ due to commutation, we observe that $g^{m}\circ f^{2m}\circ g^{m}(w_{0})$ cannot be bounded (as $w_{0}$ will either be in $BU(g)$ or $I(g)$ and $f(BU(g)\subset BU(g)$, $f(I(g)\subset I(g)$).
-Hence, $g^{m}\circ f^{2m}\circ g^{m}(w_{0})=f^{m}\circ g^{2m}\circ f^{m}(w_{0})$ in which RHS is bounded while LHS is not which is a contradiction.
-Similar arguments may be given when $w_{1}\in B$.
-Hence, $K(f\circ g)\not\supset K(f)\cap K(g)$.
-∎
-The next result provides condition for equality of bungee set, filled-in Julia set and escaping set for commuting functions.
-Theorem 2.4.
-Commuting transcendental entire functions with no finite asymptotic values have equal bungee sets, filled-in Julia sets and escaping sets.
-Proof.
-We first establish that $BU(f\circ g)\subset BU(f)\cap BU(g).$ Let $z_{0}\in BU(f\circ g)$. By definition, there are two subsequences $\{m_{k}\}$ and $\{n_{k}\}$ both tending to $\infty$ s.t.,
-$|(f\circ g)^{m_{k}}(z_{0})|=|f^{m_{k}}\circ g^{m_{k}}(z_{0})|<R\ \text{for some}\ R$
-$|(f\circ g)^{n_{k}}(z_{0})|=|f^{n_{k}}\circ g^{n_{k}}(z_{0})|\rightarrow\infty$
-So, for $\{m_{k}\}$ subsequence, $|(f\circ g)^{m_{k}}(z_{0})|=|f^{m_{k}}\circ g^{m_{k}}(z_{0})|<R$ for some $R.$ As $f$ does not have any finite asymptotic value, so $|g^{m_{k}}(z_{0})|<R_{1}$ for some $R_{1}$.
-We observe same thing for $f$ as well i.e., $|f^{m_{k}}(z_{0})|<R_{2}$ for some $R_{2}$ (because $|f^{m_{k}}\circ g^{m_{k}}(z_{0})|=|g^{m_{k}}\circ f^{m_{k}}(z_{0})|<R$ as $f,g$ are commuting).
-Now for $\{n_{k}\}$ subsequence, $|(f\circ g)^{n_{k}}(z_{0})|=|f^{n_{k}}\circ g^{n_{k}}(z_{0})|\rightarrow\infty.$ We have 3 possible cases: $|g^{n_{k}}(z_{0})|\rightarrow\infty$ or $|g^{n_{k}}(z_{0})|\rightarrow R_{3}$ for some $R_{3}$ or $g^{n_{k}}(z_{0})$ itself exhibits bungee behaviour.
-If we have the second case, it means $z_{0}\in K(g)$ (as this will imply all subsequences of $g^{j}(z_{0})$ are bounded where $j$ is some arbitrary index unrelated to $m_{k}$ and $n_{k}$).
-As a result, $g^{j}\circ f^{2j}\circ g^{j}(z_{0})$ will remain bounded (because $g^{j}\circ f^{2j}(K(g))\subset g^{j}(K(g))$ which is bounded).
-Hence, this case isn’t possible and $z_{0}\not\in K(g)$.
-On similar lines, it can be seen that $z_{0}\not\in K(f)$ where similar arguments follow after interchanging roles of $f$ and $g$ due to commutation.
-Now, when we have either of the first or the third case, it means the same thing.
-Both these cases essentially implies that $z_{0}\in BU(g)$.
-From Theorem 2.2, we have $g^{n_{1k}}\circ f^{2n_{1k}}\circ g^{n_{1k}}(z_{0})$ exhibits bungee behaviour which is as expected.
-Due to commutation, we also conclude that $z_{0}\in BU(f)$ as well.
-Hence, $BU(f\circ g)\subset BU(f)\cap BU(g)$.
-Given the preceding proof, we ask a question:
-Question: What can we say about points of the set $BU(f)\cup BU(g)\setminus BU(f)\cap BU(g),$ where will it be situated?
-Answer: As the functions $f$ and $g$ do not have any finite asymptotic values, therefore, points of the set $BU(f)\cup BU(g)\setminus BU(f)\cap BU(g)$ cannot intersect with $K(f\circ g).$ Also, $I(f\circ g)\subset I(f)\cup I(g)$ [14] and as $(BU(f)\cup BU(g))\cap(I(f)\cup I(g))=\emptyset,$ it means that $BU(f)\cup BU(g)\setminus BU(f)\cap BU(g)$ cannot intersect with $I(f\circ g)$ either.
-This implies that the only place it will be contained is $BU(f\circ g)$.
-However, if it happens, then such points must be contained in $BU(f)\cap BU(g)$ implying that no such points exist.
-This is only possible if $BU(f)=BU(g)$.
-Similar question can be asked for filled-in Julia sets as well, (as it was shown in Lemma (2.3) that $K(f\circ g)\subset K(f)\cap K(g)$).
-And the answer is: Let $w_{0}\in K(f)\cup K(g)\ \backslash\ K(f)\cap K(g)$.
-These points won’t intersect with $I(f\circ g)$ (as $I(f\circ g)\subset I(f)\cup I(g)$ and $K(f)\cup K(g)\cap I(f)\cup I(g)=\emptyset$ [14]) nor with $BU(f\circ g)$ (as $BU(f\circ g)=BU(f)=BU(g)$ and $BU(f\circ g)\cap K(f)\cup K(g)=\emptyset)$ So, they will intersect with $K(f\circ g)$ which means $w_{0}\in K(f)\cap K(g)$ which is a contradiction.
-Hence, we conclude no such points are possible.
-Therefore, $K(f\circ g)=K(f)=K(g)$ when there are no finite asymptotic values for these commuting functions.
-Hence, given the following results:
-$K(f\circ g)$ $=K(f)=K(g)$
-and
-(2.1) $BU(f\circ g)$ $=BU(f)=BU(g)$
-one can conclude that $I(f\circ g)=I(f)=I(g)$, for we know that $K(f)\cup I(f)\cup BU(f)=\mathbb{C}=K(f\circ g)\cup I(f\circ g)\cup BU(f\circ g)=K(g)\cup I(g)\cup BU(g)$ and from the previous results (2) we can conclude that $I(f\circ g)=I(f)=I(g)$.
-∎
 The following examples illustrate the results established above.
 Example 2.8.
-Consider $f(z)=1+z+e^{-z}$(Fatou map) and $g(z)=1+z+e^{-z}+2\pi i.$ It can be easily checked that $f$ and $g$ do not have any finite asymptotic values.
-One may also consider the following functions for illustration of above result:
-$f_{1}(z)=1+z+e^{z}$ which would be commuting with $g_{1}(z)=1+z+e^{z}+2\pi i$; $f_{2}(z)=z+sin(z)$ which would be commuting with $g_{2}(z)=z+sin(z)+2\pi.$
+Consider $f(z)=1+z+e^{-z}$(Fatou map) and $g(z)=1+z+e^{-z}+2\pi i.$ It can be easily checked that $f$ and $g$ satisfies Property A.
+One may also consider the following functions for illustration of above result. They also satisfy Property A.
+$f_{1}(z)=1+z+e^{z}$ which would be commuting with $g_{1}(z)=1+z+e^{z}+2\pi i$; $f_{2}(z)=z+\sin(z)$ which would be commuting with $g_{2}(z)=z+\sin(z)+2\pi.$
 3. Some results for functions of class $\mathcal{B}$
-Now, we will prove some results when these functions indeed have some finite asymptotic values, and to simplify, we will focus on the functions of the Eremenko-Lyubich class $\mathcal{B}$.
+Now, we will prove some results when the functions belong to the Eremenko-Lyubich class $\mathcal{B}$.
 These functions do not have a Baker domain; therefore, the escaping set does not intersect with the Fatou set.
-In addition, they do not have a finite intersection of the bungee set with the wandering domain because the wandering domain of such functions does not have (sub)sequences that tend to $\infty$.
-Therefore, the Julia set for functions of class $\mathcal{B}$ comprises of $RP(f),BU(f)$ and $I(f)$; that is, $J(f)=BU(f)\cup I(f)\cup RP(f)$.
-For commuting functions $f$ and $g$, we have $J(f\circ g)\supset J(f)\cup J(g)$ [10, 11].
-However, when these are commuting functions of class $\mathcal{B}$ we have, $J(f)=J(g)=J(f\circ g)$ [10].
-We also know that $g(J(f))\subset J(f)$ and $\overline{RP(f)}=J(f),\overline{I(f)}=J(f)$ [4, 6].
+In addition, the bungee set of the function does not intersect with a wandering domain in which the iterates tend to $\infty.$ However, functions in this class may have oscillating wandering domain.
+The next result gives a sufficient condition under which the bungee set is contained in the Julia set.
 Proposition 3.1.
-The bungee set for functions of class $\mathcal{B}$ does not intersect with Fatou set.
+The bungee set for functions of class $\mathcal{B}$ having no oscillatory wandering domain does not intersect with Fatou set, that is, $BU(f)\subset J(f).$
 Proof.
-[…]
-This also means that $BU(f)\subset J(f)$ only for functions of class $\mathcal{B}.$ Also, it is easy to see that $BU(f)$ is completely invariant by the complete invariance of $I(f)$ and $K(f)$.
+For a function $f,$ we know that if $BU(f)\cap F(f)\not$$\emptyset$, then such components are oscillatory wandering domains [16] and these components are contained inside $BU(f)$.
+As $F(f)$ has no oscillatory wandering domains, it means, $BU(f)$ does not intersect with $F(f)$.
+In other words, $BU(f)\subset J(f).$
+Remark 3.2.
+This also means that $BU(f)\subset J(f)$ for functions not necessarily of $\mathcal{B}$ having no oscillatory wandering domain.
+Also, it is easy to see that $BU(f)$ is completely invariant by the complete invariance of $I(f)$ and $K(f)$.
 As a result, by the minimality property of Julia set, $\overline{BU(f)}=J(f)$.
+It is also known that for a function of $\mathcal{B},J(f)=\overline{I(f)}$ [6].
 The escaping set of a function is in general, neither open, nor closed [5]. We have similar kind of result for the bungee set.
 Proposition 3.3.
-The bungee set for functions of class $\mathcal{B}$ is neither open nor closed.
+The bungee set for functions of class $\mathcal{B}$ having no oscillatory wandering domain is neither open nor closed.
 Proof.
@@
 Now, if we assume that bungee set is closed then it will be a closed completely invariant set.
-This implies that $J(f)\subset BU(f)$ (because $J(f)$ is the smallest closed completely invariant set) which is a contradiction.
-Furthermore, if bungee set is open, it will imply that it has interior to it which is also not possible (because $BU(f)\subset J(f)$ and having interior of bungee set means Julia set also has interior which is not possible for an entire function $f$ for which $F(f)\neq\emptyset$ [3]).
+This implies that $J(f)\subset BU(f)$ (because $J(f)$ is the smallest closed completely invariant set) which is a contradiction as $J(f)$ contains periodic points.
+Furthermore, if bungee set is open, it will imply that it has interior points which is also not possible (because $BU(f)\subset J(f)$ and bungee set having interior means Julia set also has interior which is not possible for an entire function $f$ for which $F(f)\neq\emptyset$ [3]).
 Hence, bungee set is neither open nor closed.
 ∎
+The following result provides equality of bungee set, filled-in Julia set and escaping set of two functions in which one function is periodic translate of the other.
+Note that the two functions are non-commuting.
 Theorem 3.4.
-Functions of class $\mathcal{B}$ which are periodic translate of each other also have equal bungee sets, escaping sets and filled-in Julia sets.
+Functions which are periodic translate of each other have equal bungee sets, escaping sets and filled-in Julia sets.
 Proof.
-We take $f(z)$ and $g(z)=f(z)+C$ where $C$ is the periodic translate of $f$ (more generally, we can take $g=f^{l}(z)+C$, for some $l\in\mathbb{N}$).
-Then $g^{n}(z)=f^{n}(z)+C$ which is almost commuting.
-Here we can see that if a point is bounded under $f^{n},$ it is also bounded under $g^{n}.$ Similarly, if a point is escaping under $f^{n},$ it is also escaping under $g^{n}$.
-Hence, $BU(f)=BU(g),I(f)=I(g),K(f)=K(g)$.
+We consider two functions $f$ and $g$ where $g$ satisfies $g(z)=f(z)+C.$ Here, $C$ is the period of $f$ (more generally, we can take $g=f^{l}(z)+C$, for some $l\in\mathbb{N}$).
+It can be easily seen that $g^{n}(z)=f^{n}(z)+C$.
+Again, it is easy to show that if a point is bounded under $f^{n},$ it is also bounded under $g^{n}.$ Similarly, if a point is escaping under $f^{n},$ it is also escaping under $g^{n}$.
+Hence, $K(f)=K(g),I(f)=I(g)$ and $BU(f)=BU(g)$.
 ∎
 Such functions can be viewed as almost commuting type with a constant shift.
-Now moving forward, we will study functions that are exactly commuting of class $\mathcal{B}$ and having finite asymptotic values.
-Lemma 3.4.
-$g(BU(f))\subset BU(f)\cup RP(f)$ for commuting functions of $\mathcal{B}.$
-Proof.
-If we recall the definition of a bungee point, there must exist a subsequence which is bounded and another subsequence which is unbounded.
-Now it is obvious that when the subsequence which is bounded $f^{m}(z_{0})\rightarrow l_{0}$ (say) and when acted upon by $g$ will also keep it bounded (since $f,g$ are entire functions).
-Let us assume that there is only one such bounded subsequence and $l$ unbounded subsequence which are as mentioned below:
-$f^{m}(z_{0})\rightarrow l_{0}$
-$f^{n_{1}}(z_{0})\rightarrow\infty$
-$f^{n_{2}}(z_{0})\rightarrow\infty$
-$\vdots$
-$f^{n_{l}}(z_{0})\rightarrow\infty$
-Now, there are two possible cases when $g$ acts on those unbounded subsequences.
-Case 1. All these subsequence gets bounded, i.e.,
-$g(f^{n_{1}}(z_{0}))\rightarrow w_{1}=f^{n_{1}}(g(z_{0}))\rightarrow w_{1}$
-$g(f^{n_{2}}(z_{0}))\rightarrow w_{2}=f^{n_{2}}(g(z_{0}))\rightarrow w_{2}$
-$\vdots$
-$g(f^{n_{l}}(z_{0}))\rightarrow w_{l}=f^{n_{l}}(g(z_{0}))\rightarrow{w_{l}}$
-If this happens, then $g(z_{0})\in$ $K(f)=F(f)\cup RP(f)$.
-As $z_{0}\in J(f)$ and we know that $g(J(f))\subset J(f)$ hence, $g(z_{0})\in RP(f)$ for the above case.
-Now, let us understand what does this mean by $g(z_{0})\in RP(f)$: It means that $f^{p}(g(z_{0}))=g(z_{0})$ for some $p\in\mathbb{N}.$ This also means $g(f^{p}(z_{0}))=g(z_{0})$.
-Because $f$ and $g$ are not invertible functions we have $f^{p}(z_{0})\not=z_{0}$.
-So, $g(z_{0})\in RP(f)\ \&\ z_{0}\in BU(f)$ can be possible.
-Case 2.
-When at least one of the subsequence still remains unbounded.
-Let that be the $k-$th subsequence i.e., $g(f^{n_{k}}(z_{0}))\rightarrow\infty$ i.e., $f^{n_{k}}(g(z_{0}))\rightarrow\infty$.
-$g(f^{m}(z_{0}))\rightarrow l_{1},$ i.e.
-$f^{m}(g(z_{0}))\rightarrow l_{1}$.
-This shows that there are at least two subsequences and in this case $g(z_{0})$ belongs to $BU(f)$.
-Therefore, $g(BU(f))$ is contained in either $BU(f)$ or $RP(f)$.
-∎
-Lemma 3.5.
-$g(I(f))\subset I(f)\cup RP(f)$ for commuting functions of $\mathcal{B}.$
-Proof.
-Suppose $z_{0}\in I(f)$ and let $\lim_{n\rightarrow\infty}g(f^{n}(z_{0}))\rightarrow l_{0}$ (say) i.e., $f^{n}(g(z_{0}))\rightarrow l_{0}.$ This means $g(z_{0})\in K(f).$ But $z_{0}\not\in K(f)$ as it belongs to $I(f)$.
-So $g(z_{0})\in RP(f)$ for the above case.
-Now, suppose $\lim_{n\rightarrow\infty}g(f^{n}(z_{0}))\rightarrow\infty$ i.e., $f^{n}(g(z_{0}))\rightarrow\infty$.
-Hence, $g$ sends $I(f)$ to either $RP(f)$ or $I(f)$.
-So, $g(I(f))\subset I(f)\cup RP(f)$.
-∎
-Proposition 3.6.
-$BU(f\circ g)\subset BU(f)\cup BU(g)$ for commuting functions of class $\mathcal{B}$.
-Proof.
-Let $z_{0}\in BU(f\circ g)\ \backslash(BU(f)\cup BU(g))$.
-This implies that $z_{0}\in K(f)\cup K(g)$ or $z_{0}\in I(f)\cup I(g)$.
-By definition of bungee set, there are subsequences $\{n_{k}\}$ and $\{m_{k}\}$ tending to $\infty$ satisfying
-$|f^{n_{k}}(g^{n_{k}}(z_{0}))|\rightarrow\infty$
-$|f^{m_{k}}(g^{m_{k}}(z_{0}))|\rightarrow l_{0}\ \ \text{for some }l_{0}.$
-First, we assume that $z_{0}\in K(g)$ which in turn implies $z_{0}\in RP(g)$ (there is a clear contradiction in the $F(g)$ case because $F(g)=F(f\circ g)$), and as a result, $f^{n}(RP(g))\subset RP(g)$ and $g^{n}(f^{n}(RP(g)))$ which will be bounded.
-As can be seen, there are no subsequences that are unbounded.
-Similarly, there are no such points in $K(f)$.
-Therefore, $z_{0}\not\in K(f)\cup K(g)$.
-Now, we assume $z_{0}\in I(g)$ which means there are two possible cases
-(3.1) $|f(g^{n}(z_{0}))|\rightarrow l\ \text{for some}\ l$
-(3.2) $|f(g^{n}(z_{0}))|\rightarrow\infty$
-where $n$ denotes an arbitrary index unrelated to $n_{k}$ and $m_{k}$.
-If we have case (3.2), then from the array of sequences argument, we see that all subsequences of $(f\circ g)^{n}(z_{0})$ tend to infinity.
-If we have case (3.1), there are further possible cases: $l\in I(f)$ or $l\in K(f)$ or $l\in BU(f)$.
-Again, if we have the first two cases, then we observe that entire composite sequence $(f\circ g)^{n}$ is (un)bounded.
-The third case is also not possible.
-This is because using the previous Lemma (3.4), we have $|f^{n_{k}}(g^{n_{k}})(z_{0})|$ or $|f^{m_{k}}(g^{m_{k}})(z_{0})|$ which equals $|g^{n_{k}}(f^{n_{k}})(z_{0})|$ or $|g^{m_{k}}(f^{m_{k}})(z_{0})|$ will be either completely bounded or unbounded.
-Therefore, $z_{0}\not\in I(g)$.
-On similar lines, $z_{0}\not\in I(f)$.
-Therefore, $z_{0}\not\in I(f)\cup I(g)$.
-Thus, we see that there are no points s.t., $z_{0}\in BU(f\circ g)\ \backslash(BU(f)\cup BU(g))$.
-∎
-Corollary 3.7.
-For commuting functions of class $\mathcal{B},(I(f)\cup I(g))\cap BU(f\circ g)=\emptyset$
-Proof.
-We can follow the similar lines of proof that is done in the preceeding Proposition (3.6).
-Let $z_{0}\in I(f)$ (say).
-Then $g(f^{n}(z_{0}))\rightarrow\infty$ or $L$.
-If $g(f^{n}(z_{0}))\rightarrow\infty,$ then by array of sequences argument, we conclude that $g^{n}\circ f^{n}(z_{0})\not\rightarrow M$ for some $M$ which means that we do not have a subsequence that remains bounded upon iteration of $f\circ g$.
-Moreover, if $g(f^{n}(z_{0}))\rightarrow L$ then $g(z_{0})\in RP(f)$ and therefore, $f^{n}(g^{n}(f^{n}(z_{0})))\rightarrow L_{1}$ (say).
-Thus, we never get a subsequence that tends to $\infty$.∞ .
-Similarly, it can be argued when $z_{0}\in I(g)$.
-∎
-Revision 3.8.
-Revised proof of $I(f\circ g)\subset I(f)\cup I(g)$ for commuting functions of $\mathcal{B}.$ (considering the existence of bungee set)
-Proof.
-The motivation for the proof comes from [14].
-Let $z_{0}\in I(f\circ g)\ \backslash(I(f)\cup I(g))$ i.e., $z_{0}\in RP(f)\cup RP(g)$ or $z_{0}\in BU(f)\cup BU(g)$.
-By definition of escaping set: $f^{n}\circ g^{n}(z_{0})=(f\circ g)^{n}(z_{0})\rightarrow\infty$.
-Assume $z_{0}\in BU(f)$, then $g(BU(f))$ goes to either $BU(f)$ or $RP(f)$ using Lemma 3.4.
-When it goes to $RP(f)$, we know $g(RP(f))$ will remain in $RP(f)$ and upon further iteration of $f\circ g$ will stay bounded i.e., $(g\circ f)^{n}(z_{0})\rightarrow L$ as $n\rightarrow\infty$ which is a contradiction.
-If it goes to $BU(f)$ then again, its the same case as discussed before when $g$ acts on it.
-In the end, it will either be an entirely bounded sequence or there would be an (un)bounded subsequence.
-The later case arises only when $g^{n}(BU(f))\subset BU(f)\ \forall n\in\mathbb{N}$ (where $n$ is some arbitrary index).
-But by definition, escaping sets must have all the subsequences tending to infinity.
-Hence, we arrive at a contradiction.
-On similar lines, it can be shown that when $z_{0}\in BU(g)$ we again arrive at a contradiction.
-So, $z_{0}\not\in BU(f)\cup BU(g)$.
-Now the case where $z_{0}\in RP(f)\cup RP(g)$.
-Assuming $z_{0}\in RP(f)$ and as discussed earlier we will have all bounded subsequences of $f^{n}\circ g^{n}(z_{0})$ which is a contradiction.
-Similar arguments can be given when $z_{0}\in RP(g).$ Therefore $z_{0}\not\in RP(f)\cup RP(g).$ Hence, $z_{0}\in I(f\circ g)\ \backslash(I(f)\cup I(g))$ is not possible.
-Thus, $I(f\circ g)\subset I(f)\cup I(g)$.
-∎
-Proposition 3.9.
-$RP(f\circ g)\supset RP(f)\cap RP(g)$ for commuting functions of class $\mathcal{B}$.
-Proof.
-$BU(f\circ g)\cup I(f\circ g)\subset BU(f)\cup BU(g)\cup I(f)\cup I(g)$
-Taking complement on both sides
-$BU(f\circ g)^{c}\cap I(f\circ g)^{c}\supset BU(f)^{c}\cap BU(g)^{c}\cap I(f)^{c}\cap I(g)^{c}$
-(3.3) $\implies RP(f\circ g)\supset RP(f)\cap RP(g)$
-Remark 3.10.
-For $f\in\mathcal{B},$ we have $J(f)=\overline{I(f)}=\overline{BU(f)}=\overline{RP(f)}=I(f)\cup BU(f)\cup RP(f).$
 4. Illustrations:
 We now illustrate $BU(f\circ g)\subset BU(f)\cup BU(g)$ with some examples. Our first example is commuting rational functions.
-Consider the functions $f(z)=z^{2}$ and $g=1/z^{2}$.
+Consider the functions $f(z)=z^{2}$ and $g(z)=1/z^{2}$.
 Observe that $f\circ g(z)=1/z^{4}$.
@@
 Our next example is motivated from [15].
-Let us take the case where $f=p(z)\cdot e^{h(z)}$ where $p(z),\ h(z)$ are polynomials and $g(z)=a\cdot f(z)$ for some $a\in\mathbb{C}\setminus\{0\}$.
-For simplicity, let us take $f(z)=z\cdot e^{a^{(k+1)}\cdot z^{k}},\ k\in\mathbb{N}$ $g(z)=a\cdot z\cdot e^{a^{(k+1)}\cdot z^{k}}$.
+Let us take the case where $f=p(z)\cdot e^{h(z)}$ where p(z),h(z)p(z),\ h(z) are polynomials and $g(z)=a\cdot f(z)$ for some $a\in\mathbb{C}\setminus\{0\}$.
+For simplicity, let us take $f(z)=z\cdot e^{a^{(k+1)}\cdot z^{k}},\ k\in\mathbb{N}$ & $g(z)=a\cdot z\cdot e^{a^{(k+1)}\cdot z^{k}}$.
 $f\circ g(z)=f(g(z))=g(z)\cdot e^{a^{(k+1)}\cdot g^{k}(z)}=aze^{a^{(k+1)}z^{k}}e^{a^{(2k+1)}z^{k}e^{ka^{(k+1)}z^{k}}}$
 $g\circ f(z)=g(f(z))=a\cdot f(z)\cdot e^{a^{k}\cdot f^{k}(z)}=aze^{a^{k}z^{k}}e^{a^{2k}z^{k}e^{ka^{k}z^{k}}}.$ Now, for $f\circ g=g\circ f,$ we need to have $a^{2k+1}=a^{k}\implies a^{k}(a^{k+1}-1)=0,$ which means $a$ is $(k+1)$-root of unity except ‘1’ itself because that’s trivial.
-Now, for the simplest case let us take: $k+1=2\implies a=\pm 1$ Since a=1 is a trivial case so let us take $a=-1$, $f(z)=ze^{-z^{2}};\ g(z)=-ze^{-z^{2}}$.
-Here as the orbit of any point under either of the function tends to $\infty,f(z)\rightarrow 0,g(z)\rightarrow 0$ and $z=0$ is a fixed point of $f(z),\ g(z)$ since $f(0)=0$ and $g(0)=0$.
+Now, for the simplest case let us take: $k+1=2\implies a=\pm 1$ Since $a=1$ is a trivial case so let us take a=1a=-1, $f(z)=ze^{-z^{2}};\ g(z)=-ze^{-z^{2}}$.
+Observe that $z=0$ is an asymptotic value of $f$ and $g$.
+It is also a fixed point of f(z),g(z)f(z),\ g(z) since $f(0)=0$ and $g(0)=0$.
 Now for the multiplier $\lambda$, $f^{\prime}(z)=e^{-z^{2}}(1-2z^{2}),f^{\prime}(0)=1$ which is an indifferent fixed point.
-And $11^{1}=$ 1 means $0\in J(f)$ or $J(g)$ (since both of them are equal).
+And $1^{1}=1$ means $0\in J(f)$ or $J(g)$ (since both of them are equal).
 As $\infty$ is a special case and belongs to $J(f)$ so, this proves that a sequence which tends to $\infty$ if it asymptotes to a finite asymptotic value which is a member of the Julia set of $f$.
-Moreover, the functions which are of the form $f=p(z)e^{h(z)}$ have a finite asymptotic value ${}^{\prime}0^{\prime}$ and it is in $J(f)$ (rationally indifferent fixed point or repelling periodic point).
 Now some illustrations for the bungee set:
-Let $f(z)=ze^{z^{2}};\ g(z)=-ze^{z^{2}}$ $\implies f^{2}=ze^{z^{2}}e^{z^{2}e^{2z^{2}}}\ \&\ g^{2}=ze^{z^{2}}e^{z^{2}e^{2z^{2}}}.$ $f\circ g=-ze^{z^{2}}e^{z^{2}e^{2z^{2}}}=g\circ f$.
+Let $f(z)=ze^{z^{2}};\ g(z)=-ze^{z^{2}}.$ This implies $f^{2}=ze^{z^{2}}e^{z^{2}e^{2z^{2}}}\ \&\ g^{2}=ze^{z^{2}}e^{z^{2}e^{2z^{2}}}.$ $f\circ g=-ze^{z^{2}}e^{z^{2}e^{2z^{2}}}=g\circ f$.
 And $(f\circ g)^{2}=g^{4}=f^{4}$(because as $(f\circ g)^{2}$ is done it does away with the negative term and rest terms are exactly same).
-So, $BU(f^{4})=BU(g^{4})=BU((f\circ g)^{2})\implies BU(f)=BU(g)=BU(f\circ g)$ (NOTE: If we have $f(z)=ze^{-z^{2}};\ g(z)=-ze^{-z^{2}}$, they also commute).
+So, $BU(f^{4})=BU(g^{4})=BU((f\circ g)^{2})$ which implies $BU(f)=BU(g)=BU(f\circ g)$ (NOTE: If we have $f(z)=ze^{-z^{2}};\ g(z)=-ze^{-z^{2}}$, they also commute).
 We may also take $f(z)=ze^{a^{(k+1)}z^{k}}$, $g(z)=aze^{a^{(k+1)}z^{k}}$ and $a^{(k+1)}=1$ for that matter).
@@
 So, $f(z)=ze^{a^{k}z^{k}}=ze^{z^{k}}$ and $g(z)=aze^{z^{k}}$.
-Now $g^{k}(z)=a^{k}\cdot f^{k}(z)=f^{k}(z)\implies BU(f)=BU(g)$.
+Now $g^{k}(z)=a^{k}f^{k}(z)=f^{k}(z)$ implies $BU(f)=BU(g)$.
 Also, $(f\circ g)^{k}(z)=a^{k}f^{2k}(z)=f^{2k}(z)$.
@@
 $J(f)=J(g)$ also holds true because we are taking $f=p(z)e^{g(z)}$ as discussed in [15].
-5. Conclusion:
-We have shown that entire functions with no finite asymptotic values which are also commuting i.e., $f\circ g=g\circ f$ satisfies:
-$K(f)=K(g),\ I(f)=I(g),\ BU(f)=BU(g)$ and therefore, $K(f\circ g)=K(f),\ BU(f\circ g)=BU(f),\ I(f\circ g)=I(f)$.
-And for the functions of class $\mathcal{B}$ which are commuting
-$RP(f\circ g)\supset RP(f)\cap RP(g)$;
-$BU(f\circ g)\subset BU(f)\cup BU(g)$;
-$I(f\circ g)\subset I(f)\cup I(g)$.
-There are some other special cases where we have equality among these sets, like when $f$ is an exponential map.
-Also, cases like $g=f+C$ where $C\in\mathbb{C}$ is the period of $f$.
-Also, when functions are from class $\mathcal{B}$, $(BU(f)\cup BU(g))\cap RP(f\circ g)\neq\emptyset$ because $g(BU(f))\subset RP(f)\cup BU(f)$.
-Similarly, $(I(f)\cup I(g))\cap RP(f\circ g)\neq\emptyset$ because $g(I(f))\subset RP(f)\cup I(f)$.
-Also, there are no points which are in $I(f)\cup I(g)$ as well as in $BU(f\circ g)$.